### v6

# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

Abstract
Virtual engines can generate dense depth maps for various synthetic scenes, making them invaluable for training depth estimation models. However, discrepancies between synthetic and real-world colors pose significant challenges for depth estimation in real-world scenes, especially in complex and uncertain environments encountered in unsupervised monocular depth estimation tasks. To address this issue, we propose Back2Color, a framework that predicts realistic colors from depth using a model trained on real-world data, thus transforming synthetic colors into their real-world counterparts. Additionally, we introduce the Syn-Real CutMix method for joint training with both real-world unsupervised and synthetic supervised depth samples, enhancing monocular depth estimation performance in real-world scenes. Furthermore, to mitigate the impact of non-rigid motions on depth estimation, we present an auto-learning uncertainty temporal-spatial fusion method (Auto-UTSF), which leverages the strengths of unsupervised learning in both temporal and spatial dimensions. We also designed VADepth, based on the Visual Attention Network, which offers lower computational complexity and higher accuracy than transformers. Our Back2Color framework achieves state-of-the-art performance on the Kitti dataset, as evidenced by improvements in performance metrics and the production of fine-grained details. This is particularly evident on more challenging datasets such as Cityscapes for unsupervised depth estimation. Code is available at https://github.com/BlueEg/back2color/.
Index Terms:
Back2Color, Syn-Real CutMix, auto-learning uncertainty, Visual Attention Depth Network
Figure 1: Performance vs. methods efficiency on KITTI at a resolution of 320 × 1024. Our method(B2C) surpasses both the lightweight Lite-mono[1] and the highly effective MonoViT[2], excelling in both efficiency and accuracy.
I Introduction
In the real world, even after prolonged evolution, humans are still limited to qualitatively analyzing the depth of specific points in their field of view to assist in their actions[3]. Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[4, 5]. Even with the refinement of advanced LiDAR systems[5], their data are inherently sparse, comprising only discrete depth points. Unsupervised depth estimation(UDE), powered by epipolar geometry[6] using purely visual approaches does not require additional annotation for depth information other than camera parameters. UDE eliminates the necessity of annotating depth labels for specific scenes. By utilizing binocular and motion disparity clues, it mimics the structure of human vision, making it efficient and cost-effective[7]. This is why many researchers and companies are actively pursuing this approach[8]. Monodepth2[9], the most renowned framework for UDE, has established itself as a classic and successful method. By reprojecting related frames(temporal and spatial) and regressing depth values. This adaptability enables Monodepth2 to obtain depth information for various scenes through adaptive fine-tuning, making it highly versatile. However, despite its strengths, Monodepth2 has several limitations. Firstly, the auto-masking for handling non-rigid moving objects, follows the rule: the pixel error after reprojection should be less than before reprojection. This pixel-level determination cannot actively detect motion[10] and cannot ensure effectiveness in smooth regions. Additionally, when applied to more complex datasets such as Cityscapes[11], it almost fails. Secondly, photometric reprojection error also fails in areas with no texture or with texture-less background objects, leading to depth prediction failure. Lastly, methods based on photometric reprojection matching lack prior knowledge of the objective structure. These issues make it challenging for UDE methods to be applied in real world.
With the continuous advancement of computer graphics and virtual engine technologies, significant progress has been made in the creation and application of virtual urban scene datasets[12, 13, 11, 14, 15]. The evolution of 3D virtual engines, such as Unity and Unreal Engine[16, 17], has played a crucial role in enhancing the realism and visualization effects of virtual urban scenes. These engines not only facilitate finer scene modeling but also provide interactivity and dynamism, making virtual urban scene data more closely resemble the real world. Synthetic datasets can provide precise qualitative data for virtual urban scenes, including depth, semantic labels, and more. Therefore, it would be desirable if various synthetic scenes and dense depth could be used to assist in the UDE of the real-world[18]. Compared to synthetic datasets, real-world depth estimation datasets[19, 11] have the following notable drawbacks: firstly, synthetic depth data is dense, whereas depth data collected by LiDAR is sparse and discrete[19]. Secondly, depth computed through stereo matching suffers from occlusions and lacks accuracy due to the short baseline[11]. Additionally, sparse depth sampling does not provide sufficient information for learning when obstacles are small or distant (resulting in small apparent areas), making it challenging for the network to accurately reconstruct their shapes.
Figure 2: Illustration of our framework. Our Back2Color training follows a two-step optimization process within each iteration. In the first step, we use Optimizer 1 to train the Depth Net for supervised Syn-Real depth estimation and the Color Net to predict real colors from depths. Initially, ① we freeze the Depth Net and Color Net to infer both real and synthetic predicted depths, using synthetic depth data to predict Syn-to-Real Color. Next, ② we apply Syn-Real CutMix to generate new syn-real colors, and ③ optimize the Depth Net with synthetic ground truth and the Color Net on the real dataset. In the second step, with Optimizer 2, ④ we train for real-world unsupervised depth estimation, incorporating sky regularization loss and auto-learning uncertainty techniques. Note that the dotted line indicates inference only.
In this paper, our Back2Color learns the mapping process in the target domain (Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains. Additionally, we introduced an improved version of CutMix[20], termed Syn-Real CutMix, to effectively combine real-world, synthesized, and syn-to-real data in both batch and image spatial dimensions. By incorporating Syn-Real CutMix into our training process, we were able to enrich the training dataset with more diverse color-to-depth knowledge extracted from synthetic scenes. Furthermore, the joint training with diverse data sources enables the model to capture a wider range of color and depth variations present in real-world scenes, thus enhancing its robustness and generalization capability. Additionally, aiming to narrow the gap between predicted depths from real-world and synthesized depth, we addressed the issue of chaotic depth clouds (5) by introducing a Background Textureless Regularization Loss, also known as Sky Regularization Loss. Through joint training, our method enables the unsupervised monocular depth estimation to leverage knowledge from synthesized data, resulting in the generation of dense and sharp depths.
In pursuit of higher prediction accuracy and a more lightweight model compared to Transformer-based methods, we introduced the Visual Attention Network[21](VAN) as the backbone and equipped it with an efficient single-path high-resolution decoder, forming the new Visual Attention Depth Net(VADepth). Our design demonstrated more effective performance than Transformers. In our in-depth experiments, we found that Auto-masking is not always effective in identifying non-rigid moving objects with the same direction and speed. Additionally, it cannot handle vehicles moving in opposite directions, leading to significant errors in depth estimation. Therefore, we introduce the automatic learning uncertainty estimation fusion method[22], effectively combining the advantages of temporal-spatial unsupervised approaches, termed Auto-learning Uncertainty Temporal-Spatial Fusion(Auto-UTSF). This method thoroughly resolves the issue of non-rigid motion in unsupervised estimation, greatly improving the accuracy of depth estimation. To summarize, the key contributions of our work include:
  • To address the issue of color and texture discrepancies between synthetic and real-world datasets, we propose Back2Color. This framework stores the color information of specific scenes (target domain) within the model, enabling the reverse prediction of color images from depth data. Additionally, we introduce Syn-Real CutMix to harmonize the joint training of real-world and synthetic data, thereby enhancing the effectiveness of unsupervised depth estimation in real-world scenes.
  • We designed an efficient and lightweight Depth Net based on the Visual Attention Network. This network boasts fewer parameters and lower computational complexity compared to transformer-based networks, while delivering superior performance.
  • To more robustly recognize non-rigid moving objects and address occlusion issues in stereo systems, we enhance the method based on Auto-Uncertainty by introducing Auto-Spatial-Temporal Uncertainty Fusion, which leverages complementary spatial and temporal advantages in the unsupervised depth estimation process.
  • Experimental results demonstrate that our method surpasses all transformer-based models, achieving state-of-the-art (SOTA) metrics while maintaining manageable computational costs. Additionally, the visual boundaries of the estimated depth images are clearer.
Figure 3: Color transformation for specific datasets: we showcase the capability of our Back2Color framework to effectively transform various synthetic datasets, including vKitti2[12], SHIFT[23], and UrbanSyn[24], into Syn2Real images. These images closely emulate the real-world styles of Kitti[25, 19] and Cityscapes[11], demonstrating high fidelity and credibility.
II Related Works
II-A Unsupervised Depth Estimation
Unsupervised depth estimation aims to predict distances from stereo views or video sequences. Initial methods like those by Godard et al.[26] introduced disparity consistency loss, while Zhou et al.[27] pioneered depth and pose estimation simultaneously. Recent advancements[28, 29, 30, 31, 32, 33, 34, 35] refine network structures for higher accuracy. Techniques like PackNet[28] and CADepth[29] enhance geometric and appearance information propagation. HR-Depth[30] and DIFFNet[31] utilize HRNet for better resolution utilization. RA-Depth[32] employs HRNet for both encoding and decoding. In natural language processing, the Transformer architecture[36] has been successful, later adapted to computer vision tasks[37, 38]. It serves as the backbone for depth estimation[39, 40, 2, 1, 41], but faces challenges like computational demands and real-time constraints. Existing methods prioritize model architecture, often overlooking crucial obstacles and targets. Additionally, depth estimation models also face the issue of overfitting on specific datasets and a lack of continuous learning capability[42].
II-B Depth Estimation with Synthetic data
Researchers have already attempted to use synthesized dense depths to guide the training of real-world depth estimation[43, 18]. Gurram et al.[18] categorized this problem as a dimension adaptation problem, where the objective is to adapt models trained on a source(synthesized) domain to perform well in a different target domain(real-world). They employed the Gradient Reverse Layer[44] to identify features from different domains during back propagation. Swami et al.[43], on the other hand, simultaneously trained on both synthesized and real-world datasets, leveraging scale information in Synthetic dataset to recover scale in unsupervised depth estimation. To bridge the gap between real-world and synthetic datasets in the image domain, Zhao et al.[45] leveraged synthetically rendered data and employed image translation techniques to narrow the synthetic-real domain gap, thereby enhancing monocular depth prediction.
II-C Style transfer in urban scenes
Research on style transfer for urban scene images has long been ongoing due to its importance in autonomous driving technology. Isola et al.[46] proposed a conditional GAN framework for generating urban scene images from semantic segmentation images of Cityscapes in 2017. However, generative models may produce random and plausible objects, disrupting the scene structure, as shown in their experiments. Atapour et al.[47] utilized style transfer and adversarial training to convert real-world color images into synthetic environment-style images. Huang et al.[48] applied GAN-based domain transfer for data augmentation in urban scene weather-time conversion. To minimize the gap between source and target domains, Gomez et al.[49] aligned colors across different domains in the Lab color space. Zheng et al.[50] used GANs to convert synthetic colors to real-world colors.
III Methodology
In this section, we introduce the Back2Color framework, which masters the depth-to-color mapping within the target domain to train the Color Net (III-A1). Back2Color is enhanced by Real-Syn CutMix (III-A2) and a strategy for handling textureless background objects (III-A3). We then discuss the VADepth, based on the Visual Attention Network, designed for high performance and lightweight architecture (III-B). Lastly, we describe the Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) (III-C) for addressing non-rigid motion challenges. In this paper, Depth and Disparity are equivalent.
Figure 4: Syn-Real CutMix: specific steps involved in our Syn-Real CutMix method for spatially combining colors: (a) Randomly select a rectangular region, denoted as M. (b) Replace the corresponding block in image $I_{2}$ with the block from image $I_{1}$. (c) Resulting in the formation of new paired data $(\tilde{I},D)$.
III-A Back to the Color
III-A1 Transform depth to specific colors
the significant gap between synthetic and real-world color images arises from the inherent complexity of the real world. Colors in real-world scenes are influenced by various factors, including lighting intensity, surface material of obstacles, and smoothness, and can also be significantly degraded[51]. Each of these factors is intricate and highly variable, making them challenging to simulate realistically using typical 3D engines. As a result, discerning the origin of an image intuitively is easy, both for humans and deep learning alike. Consequently, directly combining real-world and synthetic datasets for training essentially involves dealing with the integration issues of two distinct domains. Our Back2Color approach learns the mapping relationship in the target domain(Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains.
Objectively speaking, predicting color from 3D structural information constitutes a one-to-many mapping, a non-standard conditional generative tasks, where the same 3D structure can correspond to a multitude of different color images. However, under specific conditions, this seemingly impossible task becomes feasible: Assuming that the color images in the target domain exhibit relatively high correlation. The colors in real word datasets like Cityscapes and Kitti are notably different, as shown in Fig.3. Therefore, in order to bridge the substantial gap between synthetic $\mathcal{S}\in${vKitti2[12](Vk),SHIFT[23](S), UrbanSyn[24]} and real-world $\mathcal{X}\in$Kitti[25, 19](K),Cityscapes[11](Cs)}datasets, we can learn the depth-to-color mapping in the target domain (real world color) and apply it to the source domain (synthetic color), thereby converting the source domain color to match those in the target domain. And then, we made the following attempts, train a Color Net $\mathcal{F}_{d2c}$ to predict color images $\tilde{\mathbf{I}}^{\mathcal{X}}$ from predicted depth images $\hat{\mathbf{D}}^{\mathcal{X}}$:
$\hat{\textbf{I}}^{\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\textbf{D}}^{\mathcal{X}}|\boldsymbol{\theta}^{\mathcal{X}}),$ (1)
and utilizing the input of the Depth Net $\textbf{I}^{\mathcal{X}}$ for Supervision:
$\mathcal{L}^{d2c}=pe(\hat{\textbf{I}}^{\mathcal{X}},\textbf{I}^{\mathcal{X}}),$ (2)
Figure 5: Chaotic depth clouds: The green pixels can be matched across relevant frames based on the features of foreground obstacle, while the red pixels lacks effective features for matching, resulting in a cloud-like region in the depth image.
where $\boldsymbol{\theta}^{\mathcal{X}}$ represents the weights that store specific dataset color mapping information, such as Kitti and Cityscapes. Fortunately, we found that this prediction is feasible, neural networks are able to capture color information under specific target scenes. Next, we need to investigate whether we can employ $\boldsymbol{\theta}^{\mathcal{X}}$ to convert predicted $\hat{\textbf{D}}^{\mathcal{S}}$ in synthesized datasets into color that resemble the Kitti($\mathcal{X}$) color style: Using the specific Color Net to colorize predicted synthetic depth, as shown in Fig.3:
$\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\textbf{D}}^{\mathcal{S}}|\boldsymbol{\theta}^{\mathcal{X}}).$ (3)
Figure 6: The proposed depth estimation module based on the Visual Attention Network. The encoder is composed of an ImageNet pretrained Visual Attention Network (VAN), generating multi-scale features. The VAN stages consist of the Large Kernel Attention (LKA) module, which is a pure CNN structure. This encoder balances performance and efficiency. The decoder features an HRNet structure, where each stage fuses multi-scales of features upward.
Our experiments have shown that by referencing color information from the target domain, we can learn its depth-to-color mapping and accurately predict color images in the source domain that closely resemble those in the target domain.
III-A2 Syn-Real CutMix
CutMix[20] replaces removed pixels with patches from other images rather than simply removing them. This strategy eliminates uninformative pixels during training, enhancing efficiency, while maintaining the advantages of regional dropout to focus on non-discriminative obstacle parts. Additionally, introducing extra patches improves the model’s localization ability, requiring identification of obstacles from partial views. In our work, we draw inspiration from the CutMix method to achieve different objectives and effects. We aim to dynamically fuse Syn2Real Colors and Synthetic Colors within a single batch for supervised training, a technique we have named Syn-Real CutMix.
The goal of CutMix in our work is to generate a new training pair $(\tilde{\textbf{I}},\textbf{D})$. Unlike CutMix[20], our samples$(\textbf{I}^{\mathcal{S}},\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}})$ are not two unrelated samples; instead, they correspond to different colored images of the same depth label $\textbf{D}^{\mathcal{S}}$. Therefore, we only merge color images without combining new labels. To increase the randomness of training samples. As shown in Fig.4, given image samples $(\textbf{I}^{\mathcal{S}},\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}})$ as $(\textbf{I}_{1},\textbf{I}_{2})$, we firstly randomly swap $(\textbf{I}_{1},\textbf{I}_{2})$. Then, randomly crop a block from $\textbf{I}_{1}$ with a height and width greater than half of the original image, and replace it with the corresponding position in $\textbf{I}_{2}$:
$\tilde{\textbf{I}}=\textbf{M}\odot\textbf{I}_{1}+(1-\textbf{M})\odot\textbf{I}_{2},$ (4)
where $M_{i}\in\{0,1\}$ denotes a binary mask indicating where to drop out and fill in, and $\odot$ is element-wise multiplication.
Now, we combine Syn-Real enhanced images $\tilde{\textbf{I}}$ and Real world target image I into one batch for training. Then, we can train monocular depth estimation networks simultaneously on real-world, synthetic, and Syn-to-Real images within a batch. Spatial randomization also encourages the network to focus more on the image structure rather than its specific colors.
III-A3 Background textureless regularization loss
when analyzing and comparing unsupervised monocular depth and synthesized depth, we found not only significant gap in color images but also substantial gaps in the depth outputs generated by the model. In Fig.5, when the background of obstacle is textureless regions, these areas lack features for matching in sequences or stereo, making it challenging to accurately determine their depth. Consequently, these regions are prone to arbitrary depth assignments within the textureless area, leading to the formation of cloud-like regions in the depth map, we called them chaotic depth clouds. The main drawback of such cloud-like regions lies in their ability to significantly impact the overall distribution of depth values in the depth map, resulting in notable biases when calculating median variance during evaluation processes.
For chaotic depth clouds, they appear mostly within regions with the sky as background, such as trees, road signs, and buildings. Fortunately, the sky’s depth itself has a strong prior, that it is infinitely far away. Therefore, we predict the mask of the sky and constrain its value on the disparity map estimation by the network. This ensures that the network separates background and foreground obstacles:
$\mathcal{L}^{sky}=\frac{\sum|\textbf{M}_{sky}\times\hat{\textbf{D}}|}{\sum\textbf{M}_{sky}},$ (5)
where $\textbf{M}_{sky}$ is inferred from a pre-trained semantic segmentation network[52]. We also have an intuitive prior that the sky should only exist above in the view, so we added a decay coefficient $\textbf{G}\in\mathbb{R}^{W\times H}$ that decreases with the image height:
$\textbf{G}[..,j]=(H-j)/H,$ (6)
where $H$ is the height of $\textbf{D}_{t}$, so our final sky regularization loss function is:
$\mathcal{L}^{sky}=\frac{\sum\textbf{G}\times|\textbf{M}_{sky}\times\hat{\textbf{D}}|}{\sum\textbf{M}_{sky}}$ (7)
The proposed sky regularization loss effectively eliminates hazy depth while enhancing the recovery of delicate objects.
Figure 7: When network capabilities become stronger, and the number of non-rigid objects in the dataset increases, the shortcomings[22] of Auto-masking and Minimum Reprojection Error lead to the reappearance of errors.
III-B Visual Attention Depth Net
Every high-performing depth estimation network requires an effective and robust backbone to extract hierarchical features from natural images. Commonly used backbones include ResNet[53] (Monodepth2[9], HRDepth[30], CADepth[29]), HRNet[54] (DIFFNet[31], RA-Depth[32]), and Transformer-based backbones like ViT[37], MPViT[41](TSUDepth[22], MonoViT[2]). With the evolution from ResNet to HRNet and then to Transformer-based approaches, the performance of depth estimation continues to improve, but accompanied by increasing computational complexity and params. We consistently believe that depth estimation tasks, especially those based on street scenes, need to consider the requirements of their usage scenarios (real-time, lightweight). Hence, we aim to find a more efficient and effective backbone than Transformer methods.
Research on balancing the performance and efficiency of CNN and Transformer models has been ongoing, and we have discovered the VAN[21] (Visual-Attention-Network) model. The VAN[21] model employs large kernel convolutions instead of self-attention mechanisms to capture long-range dependencies. As shown in Fig.6, It decomposes the large kernel convolution into a spatial local convolution (depth-wise convolution), a spatial long-range convolution (depth-wise dilation convolution), and a channel convolution (1×1 convolution) to reduce computational complexity. This design not only avoids treating 2D images as 1D sequences but also achieves stronger accuracy on ImageNet while reducing computational complexity and params.
Due to its outstanding performance and efficiency, we have introduced the VAN model into unsupervised monocular depth estimation tasks, and adapt to some fine-tuning. The most significant change is that, to reduce computational complexity, VAN downsamples the input resolution to $2^{-2}$ during the first stem stage. While this downsampling is acceptable for classification tasks, it poses challenges for high-resolution recovery in image regression tasks. Therefore, without altering the original VAN structure, we have added an additional stem head to acquire features at $2^{-1}$ resolution, as shown in Fig.6. After extracting features from various VAN stages, we equipped them with a high-resolution depth decoder[32] to output the final depth. During implementation, we observed no difference between outputting a single-size depth image and outputting multi-scale depth images. Therefore, our network only outputs depth maps with the same size as the input.
Figure 8: Temporal reprojection involves reprojecting the previous and subsequent frames to the target frame, while spatial reprojection involves reprojecting the left or right frame to the target frame.
III-C Auto-learning uncertainty temporal-spatial fusion
When testing our designed VADepth model, we found that the Auto-masking[9] method almost failed, as shown in Fig.7. This issue is more severe in the Cityscapes[11]. Upon inspecting the output of Auto-masking, we discovered that it couldn’t perfectly shield non-rigid objects. We attribute this to the model’s robustness, the limitations of the Auto-masking method, and the complexity of the Cityscapes[11, 55].
Therefore, we need to address the issue of non-rigid objects to further enhance performance. In temporal sequential frames, there are freely moving targets, such as pedestrians and vehicles(as show in Fig.8). These do not conform to the rigid scene transformation assumptions based on depth and camera pose. The solution provided by Auto-Masking[9] is to compare the pixels before and after reprojection. If the reprojection error after reprojection is worse than without reprojection, it indicates that the pixel belongs to a non-rigid object.
$\mu=[\min_{c}pe(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})<\min_{c}pe(\textbf{I}_{t},\textbf{I}_{c})].$ (8)
However, the network can still rationalize the depth of non-rigid moving objects by pulling forward targets further away and pushing backward closer targets, making them consistent with the camera extrinsics(Pose), as show in Fig.7. In fact, non-rigid moving objects only exist in temporally related frames, while in spatially related frames (stereo settings), since they are captured at the same moment, this issue does not arise. Therefore, introducing stereo supervision into the system is the most direct and effective method, rather than merely excluding masks for non-rigid objects[55, 9].
Figure 9: Auto-Spatio-Temporal Uncertainty Fusion: The Uncertainty Generation Network utilizes a Pose Encoder and establishes an Uncertainty Decoder. It inputs the $\textbf{I}_{\{-1,+1,0,s\}}$ and the target frame $\textbf{I}_{0}$ separately to output the uncertainty map $\hat{\textbf{U}}$. Then, it generates the weight map U for different error maps through a “1-Softmax” operation.
But when comparing temporal sequence reprojection with spatial reprojection, each has its own advantages and disadvantages. First, the translation of pose between temporally related frames are larger, whereas in spatial reprojection (stereo), the baseline is fixed and small. This makes temporal-based transformations better suited for estimating distant targets. Then, temporally related frames typically include three frames: the current frame and the preceding and following frames, allowing information from the preceding and following frames to fill in occluded areas in the target frame. In spatial reprojection (stereo), regardless of whether the target frame is set as the left or right eye, there is no symmetric completion of occluded areas. Thus, we seek a method that can effectively combine the advantages of both temporal and spatial reprojections to predict better depth maps.
We referenced prior studies on Uncertainty Learning[56, 57, 22]. And in the work by Zhu et al.[22], a method was introduced to address occlusion problems by automatically predicting uncertainty and reweighting the reprojection in the reprojection process of spatially adjacent frames. Building on this idea, we aim to take it a step further by introducing a method to predict uncertainty during both temporal and spatial reprojections. This approach seeks to leverage the strengths of both temporal and spatial domains effectively.
To combine photometric loss across different frames, Zhu et al.[22] specify $\sum_{c}(1-\hat{u}_{i,c})=1,c\in\{t-1,t+1\}$. Hence, they combine the output uncertainty map $\hat{\textbf{U}}_{c}(\hat{\textbf{U}}_{c}\geq 0)$ with softmax: $\{\textbf{U}_{t-1},\textbf{U}_{t+1}\}=1-Softmax\{\hat{\textbf{U}}_{t-1},\hat{\textbf{U}}_{t+1}\}$, to ensure that $\sum_{c}u_{i,c}=1$. The final joint optimization function is written as:
$\begin{split}\mathcal{L}&=\frac{1}{N}\sum^{N}\sum_{c}u_{i,c}pe_{i}(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})\\ &\quad+\alpha\frac{1}{N}\sum^{N}\sum_{c}\hat{u}_{i,c},\quad\sum_{c}u_{i,c}=1.\end{split}$ (9)
Using this formula, the network will automatically learn the uncertain regions(areas where the loss cannot be reduced through optimization) within the loss, forming the final weights for multi-frame loss through comparison. In the temporal multi-frame reprojection loss combination, high-uncertainty regions are the frames’ occluded areas revealed due to reprojection. Meanwhile, in the spatial and temporal reprojection loss combination, high-uncertainty regions in the temporal reprojection loss map are the non-rigid motion areas that can be accurately reprojected in the spatial.
$\begin{split}\mathcal{L}^{uD}=\frac{1}{N}\sum^{N}\Bigg{(}u_{i,\tau}\sum_{c}u_{i,c}pe_{i}(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})+u_{i,s}pe_{i}(\textbf{I}_{t},\textbf{I}_{s\rightarrow t})\Bigg{)}\\ +\alpha\frac{1}{N}\sum^{N}(\sum_{c}\hat{u}_{i,c}+\sum_{\sigma}\hat{u}_{i,\sigma}),\quad c\in\{-1,+1\},\quad\sigma\in\{\tau,s\}\end{split}$ (10)
where $\tau$ and $s$ represent the temporal domain and the spatial domain, respectively, with $\alpha$ set to $1e^{-2}$. $u_{i,\tau}$ and $u_{i,s}$ represent the temporal-domain uncertainty and the spatial-domain uncertainty, $\sum_{c}u_{i,c}=1,\sum_{\sigma}u_{i,\sigma}=1$, $\{\textbf{U}_{\tau},\textbf{U}_{s}\}=1-Softmax\{\hat{\textbf{U}}_{\tau},\hat{\textbf{U}}_{s}\}$. $s$ represents the source frame in a stereo setup. Since the Kitti dataset randomly selects either the left or right camera image as the target frame, the other frame becomes the source frame.
Figure 10: Temporal Uncertainty for non-rigid objects: We can identify non-rigid objects by predicting the uncertainty during the training process, thereby eliminating their impact in the trainning. The fine-grained recognition can even identify very small non-rigid targets.
Regarding how to obtain temporal and spatial uncertainty maps, we consider referencing the method used to generate temporal uncertainty maps by inputting related frames containing uncertainty information. For spatial reprojection, we can input $\textbf{I}_{s}$ and $\textbf{I}_{0}$, but there is no single related frame that uniquely represents temporal reprojection. We attempt to input $\textbf{I}_{0}$ and $\textbf{I}_{0}$ as representatives. By comparing the output uncertainty maps when calculating the loss, we can form reasonable uncertainty maps, as shown in Fig. 9. Applying uncertainty learning in temporal reprojection can address occlusion and field contraction issues. We improved this method to fuse temporal and spatial reprojection errors. Since non-rigid objects are rigid in the spatial domain, the temporal domain generates significant uncertainty(Fig.10). Additionally, high uncertainty still arises from occluded regions during the spatial domain reprojection process. By fusing the uncertainties from both the temporal and spatial domains, we can ignore high-uncertainty areas in each domain when calculating the loss, leveraging the strengths of each domain to train better depths.
III-D Training Strategy
III-D1 Synthetic supervised depth loss
Since unsupervised depth estimation models based on video sequences cannot recover scale information, we adopt the affine-invariant mean absolute error loss used by MiDas[58] and Depth Anything[59] for joint training with synthesized depth, which has deterministic values varying with different camera parameters:
$\mathcal{L}^{sD}=\frac{1}{HW}\sum_{i}\ |d^{*}_{i}-\hat{d}^{*}_{i}|,$ (11)
where $d^{*}_{i}$ and $\hat{d^{*}_{i}}$ are the scaled and shifted pixels of predicted $\hat{\textbf{D}}$ and synthetic goundtruth D.
$d^{*}_{i}=\frac{d_{i}-median(\textbf{D})}{s(\textbf{D})},s(\textbf{D})=\frac{1}{HW}\sum_{i}|d_{i}-median(\textbf{D})|.$ (12)
Figure 11: The synthetic datasets: In vKitti2, depth cannot penetrate transparent entities such as glass windows, whereas in the SHIFT dataset, depth can penetrate transparent entities. SHIFT and UrbanSyn includes simulated pedestrians. And UrbanSyn is the synthetic dataset that most closely resembles real world.
Figure 12: The MixSyn dataset contains more samples than the real-world unsupervised datasets, requiring it to be divided into multiple epochs. While the MixSyn dataset is trained once, the real-world unsupervised datasets may be trained multiple times.
III-D2 Optimization Strategy
Swami et al.[43] highlighted that multiple networks and optimization losses in unsupervised domain adaptation methods often compete against each other and fail to contribute effectively to the optimization of the depth estimation. This lack of cooperation among different components makes it challenging to achieve best results. So here, we have set two optimizers $\{\mathcal{O}_{d2c},\mathcal{O}_{d2c}\}$ to separately train the Depth Net and the Color net, as shown in Fig.2.
Optimizer 1-$\mathcal{O}_{c2d}$: Supervised training of the Depth Net on the Syn-Real samples and the Color Net on the Real world unsupervised depth dataset for predicting color from depth, $\gamma=1e^{-3}$.
$\mathcal{L}^{\mathcal{O}_{d2c}}=\mathcal{L}^{d2c}+\mathcal{L}^{sD}+\gamma\mathcal{L}^{sm}.$ (13)
Optimizer 2-$\mathcal{O}_{c2d}$: Unsupervised training of the Depth Net on the Real world and applyed the background textureless regularization loss , $\beta=1e^{-2},\gamma=1e^{-3}$.
$\mathcal{L}^{\mathcal{O}_{c2d}}=\mathcal{L}^{uD}+\beta\mathcal{L}^{sm}+\gamma\mathcal{L}^{sky}.$ (14)
Figure 13: Visual of Depth Estimation on Kitti: Our method yields clearer depth estimates with higher correlation to corresponding colors, sharper boundaries, and fewer areas with chaotic depth clouds.
IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.I, Tab.II, and Tab.IV). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.V), perform comparisons at higher resolutions (Tab.III), and benchmark against the lightest and most effective methods (Tab.VIII). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 13, Fig. 15, and Fig. 16).
IV-A Datasets
We utilized two unsupervised monocular depth estimation datasets: KITTI[25] and Cityscapes[11], along with synthetic datasets vKitti2[12], SHIFT[23], and UrbanSyn[24]. KITTI contains 39,810 training triplets, 4,424 for validation, and 697 for evaluation, featuring diverse scenes with rich colors and occasional non-rigid objects. Cityscapes, designed for autonomous driving, provides 69,731 monocular triplets from the LeftImg8bit Sequence for training, mainly urban scenes with non-rigid pedestrians and vehicles, often in darker tones. The baseline for KITTI stereo data is 54 cm, while for Cityscapes it is 22 cm. vKitti2, an upgraded version of Virtual KITTI, uses the Unity game engine for a more photorealistic environment and provides 21,260 frames for training. SHIFT features various weather conditions and times, offers dense depth information with 1 mm resolution, and includes simulated pedestrians. We selected 153,000 frames and synthetic depths from its front-view cameras. UrbanSyn, designed to address the synth-to-real domain gap, provides 7,539 images with depth labels, offering a highly realistic synthetic dataset for various computer vision tasks.
To leverage the variety of multiple synthetic datasets, we normalized and shuffled the specifications (size, depth resolution, etc.) of the vKitti2, SHIFT, and UrbanSyn datasets to create a unified dataset called MixSyn. The MixSyn dataset contains significantly more samples compared to KITTI and Cityscapes. Therefore, during the training process, MixSyn is trained once, while KITTI or Cityscapes is trained for multiple epochs, as shown in Fig.12. We demonstrate the advantages of using MixSyn in Tab. VII.
TABLE I: Experiments on the Kitti Eigen Split using only the Monocular Sequence dataset for training.
Method Sup. Size Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
Monodepth2[9] m 192x640 0.115 0.903 4.863 0.193 0.877 0.959 0.981
CADepth[29] m 192x640 0.105 0.769 4.535 0.181 0.892 0.964 0.983
DIFFNet[31] m 192x640 0.102 0.749 4.445 0.179 0.897 0.965 0.983
MonoFormer[60] m 192x640 0.106 0.839 4.627 0.183 0.889 0.962 0.983
Lite-Mono[1] m 192x640 0.101 0.729 4.454 0.178 0.897 0.965 0.983
MonoViT[2] m 192x640 0.099 0.708 4.372 0.175 0.900 0.967 0.984
Back2Color m 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
DIFFNet[31] m 320x1024 0.097 0.722 4.345 0.174 0.907 0.967 0.984
Lite-[1] m 320x1024 0.097 0.710 4.309 0.174 0.905 0.967 0.984
MonoViT[2] m 320x1024 0.096 0.714 4.292 0.172 0.908 0.968 0.984
Back2Color m 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
IV-B Implementation Details
Our training is implemented by Pytorch[61] with NVIDIA RTX 3090Ti(24GB). We have used ADAM[62] as the optimizer. We have set the learning rate to $1\times 10^{-4}$, the same as in Monodepth2[9], and the learning rate will decay to $1\times 10^{-5}$ after 15 epochs. The other parameters are the same as $(\beta_{1}$, $\beta_{2})$ = $(0.9,0.999)$. We set the SSIM weight to $0.85$ and the smoothness weight to $\alpha=1e-3$. Furthermore, we utilized a pretrained segmentation network[52] trained on the Cityscapes semantic segmentation dataset. We evaluate the performance using depth estimation metrics[25, 63]: rmse(log), rmse, Sq.relative, Abs relative, and Accuracy.
TABLE II: Experiments on the Kitti Eigen Split using only the Monocular Sequence and Stereo datasets for training.
Method Sup. Size Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
Monodepth2[9] ms 192x640 0.106 0.818 4.750 0.196 0.874 0.957 0.979
HR-Depth[30] ms 192x640 0.107 0.785 4.612 0.185 0.887 0.962 0.982
CADepth[29] ms 192x640 0.102 0.752 4.504 0.181 0.894 0.964 0.983
DIFFNet[31] ms 192x640 0.101 0.749 4.445 0.179 0.898 0.965 0.983
MonoViT[2] ms 192x640 0.098 0.683 4.333 0.174 0.904 0.967 0.984
Back2Color ms 192x640 0.092 0.685 4.229 0.170 0.911 0.968 0.984
Monodepth2[9] ms 320x1024 0.106 0.806 4.630 0.193 0.876 0.958 0.980
HR-Depth[30] ms 320x1024 0.101 0.716 4.395 0.179 0.894 0.966 0.983
CADepth[29] ms 320x1024 0.096 0.694 4.264 0.173 0.908 0.968 0.984
DIFFNet[31] ms 320x1024 0.094 0.678 4.250 0.172 0.911 0.968 0.984
MonoViT[2] ms 320x1024 0.093 0.671 4.202 0.169 0.912 0.969 0.984
Back2Color ms 320x1024 0.087 0.644 4.079 0.166 0.920 0.970 0.985
TABLE III: Experiments on the Kitti Eigen Split at 384 $\times$ 1280 resolution.
Method Sup. Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
PackNet-SfM[28] m 0.107 0.802 4.538 0.186 0.889 0.962 0.981
SGDepth[64] m 0.107 0.768 4.468 0.186 0.891 0.963 0.982
MonoViT[2] m 0.094 0.682 4.200 0.170 0.912 0.969 0.984
PlaneDepth[65] ms 0.090 0.584 4.130 0.182 0.896 0.962 0.981
Back2Color ms 0.086 0.629 4.054 0.165 0.920 0.970 0.985
IV-C Monocular Depth Estimation Performance
Our recommended experimental setup is to use the VAN2 model (balancing efficiency and performance), utilize the MixSyn synthetic dataset, and employ spatiotemporal supervision (ms). If there are no annotations, the results should be based on this setup. The evaluation results are summarized in Tab.I and Tab.II with Kitti Eigen Split[25] and Tab.IV with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.13 and Fig.15. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[1, 2] with larger params and more complex structures(Tab.VIII).
Figure 14: Temporal-Spatial Uncertainty Fusion: White represents higher reliability in spatial reprojection, while black represents higher reliability in temporal reprojection, as shown in the legend in the bottom right corner. We can observe that non-rigid moving objects in the image tend to reproject more in space (white). Meanwhile, occluded areas and distant regions tend to reproject more in the temporal sequence (black), and the remaining areas show no significant tendency (gray).
TABLE IV: Experiments on the Cityscapes benchmark with 192$\times$640. Methods* requiring semantic information and Back2Color requiring stereo set.
Method Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
Struct2Depth*[66] 0.145 1.737 7.280 0.205 0.813 0.942 0.976
GLNet[67] 0.129 1.044 5.361 0.212 0.843 0.938 0.976
Gordon et al.*[68] 0.127 1.330 6.960 0.195 0.830 0.947 0.981
Li et al.[69] 0.119 1.290 6.980 0.190 0.846 0.952 0.982
Lee et al.*[70] 0.111 1.158 6.437 0.182 0.868 0.961 0.983
RM-Depth[55] 0.090 0.825 5.503 0.143 0.913 0.980 0.993
Back2color* 0.076 0.784 5.159 0.127 0.937 0.984 0.994
TABLE V: Experiments with methods using the Synthetic dataset (vKitti2) at 192×640 resolution and training with monocular sequential only.
Method Size Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
MonoDEVSNet[18] 192x640 0.102 0.685 4.303 0.178 0.894 0.966 0.984
Back2Color 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
Swami et al.[43] 320x1024 0.103 0.654 4.300 0.178 0.891 0.966 0.984
Back2Color 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
IV-D Ablation Study
As illustrated in Tab. VI, we conducted experiments on our Back2Color framework by replacing different model structures to demonstrate the effectiveness of our method. While the Depth Net learned to predict dense depth from synthetic data, it also disrupted the model’s latent memory in unsupervised learning and learned to generate sharp boundaries of obstacles. Based on the metrics in Tab. VI, there is an improvement across all metrics, particularly the $a1$ and $rmse\_log$ metrics representing detail accuracy. At the same time, we also observed that there was no significant difference between only using SHIFT and vKitti2. In Tab. VI, we also tested training VADepth from scratch without ImageNet pre-training. The results demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation.
Then we compared our method with other noval approaches also joint training with synthetic datasets. Our improvement remains significant in comparison, as shown in Tab.V.
Figure 15: The visual results on Cityscapes: Our method yields clearer depth estimates with higher correlation to corresponding colors, sharper boundaries, and fewer areas with chaotic depth clouds.
TABLE VI: Experiments on the Kitti Eigen Split at 192×640 resolution with different models for Back2Color.VAN* means not apply ImageNet pretrained checkpoints.
Method Model Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
Monodepth2[9] Resnet34 0.106 0.818 4.750 0.196 0.874 0.957 0.979
Back2Color Resnet34 0.104 0.808 4.617 0.186 0.883 0.962 0.982
HRDepth[30] HRDepth 0.107 0.785 4.612 0.185 0.887 0.962 0.982
Back2Color HRDepth 0.105 0.745 4.469 0.182 0.887 0.964 0.984
DIFFNet[31] DIFFNet 0.101 0.749 4.445 0.179 0.898 0.965 0.983
Back2Color DIFFNet 0.094 0.660 4.254 0.172 0.904 0.967 0.984
Back2Color VAN2* 0.123 0.997 4.996 0.202 0.858 0.953 0.979
Back2Color VAN2 0.092 0.685 4.229 0.170 0.911 0.968 0.984
Additionally, we used various synthetic datasets such as vKitti2, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Tab.VII. The results consistently achieved excellent improvement, demonstrating the strong adaptability of our method. To demonstrate that the improvement in our Back2Color method is not due to the increased training samples (Synthetic Datasets) but rather due to the diverse priors brought by various synthetic datasets, we conducted experiments using 20% and 50% synthetic samples (Rows 6, 7, and Rows 9, 10) while ensuring the total number of dataset samples remains the same as the KITTI dataset (39,810). The results are consistent with those obtained using the full dataset, proving that the enhancement of our method is not due to the increase in training samples. Additionally, when using the MixSyn mixed synthetic dataset, some metrics (rmse_log and a1) show significant improvement compared to using a single dataset, demonstrating that our method benefits from learning the variety of multiple synthetic datasets.
Finally, to demonstrate the exceptional efficiency of our model, we compared it with the current best-performing MonoViT[2] and the lightweight Lite-mono[1]. Our VAN0 model is not only lighter than Lite-mono but also delivers superior performance. Additionally, it achieves competitive results with only one-sixth of the parameters of MonoViT. Meanwhile, our VAN2 model, which has a parameter count similar to MonoViT, shows a significant performance improvement, making it the current state-of-the-art (SOTA) model.
TABLE VII: Ablation Study on the impact of Various Synthetic Datasets.
Size Syn Number of Samples Error Metric Accuracy Metric
Kitti vKitti2 SHIFT UrbanSyn abs_rel sq_rel rmse rmse_log a1 a2 a3
192x640 vKitti2 39810 21260 0 0 0.094 0.694 4.296 0.173 0.907 0.968 0.984
192x640 SHIFT 39810 0 153000 0 0.092 0.676 4.238 0.171 0.909 0.969 0.984
192x640 Urban 39810 0 0 7539 0.093 0.644 4.198 0.172 0.907 0.968 0.984
192x640 MixSyn 39810 21260 153000 7539 0.092 0.685 4.229 0.170 0.911 0.968 0.984
320x1024 vKitti2 39810 21260 0 0 0.090 0.677 4.173 0.169 0.914 0.969 0.984
320x1024 vKitti2 31848 7962 0 0 0.090 0.675 4.143 0.168 0.917 0.970 0.985
320x1024 vKitti2 19905 19905 0 0 0.091 0.684 4.192 0.169 0.913 0.969 0.985
320x1024 SHIFT 39810 0 153000 0 0.089 0.651 4.158 0.167 0.916 0.970 0.985
320x1024 SHIFT 31848 0 7962 0 0.090 0.684 4.158 0.168 0.916 0.970 0.984
320x1024 SHIFT 19905 0 19905 0 0.090 0.680 4.144 0.167 0.915 0.970 0.985
320x1024 Urban 39810 0 0 7539 0.090 0.654 4.084 0.167 0.915 0.970 0.985
320x1024 MixSyn 39810 21260 153000 7539 0.087 0.644 4.079 0.166 0.920 0.970 0.985
Figure 16: The visual results of joint training with different synthetic datasets. The model trained with SHIFT jointly learns more features of pedestrians while treating glass as transparent entities.
TABLE VIII: Ablation study on models with params and computational complexities(GFLOPs).
Method Model Size Params(M) Flops(G) Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
MonoViT[2] small 320x1024 33.63 159.14 0.093 0.671 4.202 0.169 0.912 0.969 0.985
Lite-mono[1] 8M 320x1024 8.76 29.88 0.097 0.710 4.309 0.174 0.905 0.967 0.984
Back2Color VAN0 320x1024 5.17 17.64 0.094 0.706 4.257 0.171 0.910 0.969 0.984
Back2Color VAN1 320x1024 18.61 51.48 0.090 0.664 4.139 0.168 0.915 0.970 0.984
Back2Color VAN2 320x1024 31.32 76.01 0.087 0.644 4.079 0.166 0.920 0.970 0.985
Back2Color VAN3 320x1024 49.5 114.63 0.087 0.653 4.068 0.165 0.919 0.971 0.985
Back2Color VAN0 384x1280 5.17 17.64 0.093 0.676 4.169 0.171 0.911 0.969 0.984
Back2Color VAN1 384x1280 18.61 51.48 0.090 0.664 4.102 0.168 0.917 0.970 0.984
Back2Color VAN2 384x1280 31.32 76.01 0.086 0.629 4.054 0.165 0.920 0.970 0.985
Back2Color VAN3 384x1280 49.5 114.63 0.087 0.652 4.062 0.165 0.920 0.970 0.984
IV-E Visualization Analysis
The impact of synthetic datasets can be analyzed by comparing the joint training with different synthetic datasets, as shown in the Fig.16.
Firstly, it’s undeniable that our method, through joint training with synthetic datasets, achieves clearer boundaries and more detailed depth estimations compared to training solely with unsupervised depth estimation, as depicted in Fig. 13. The model successfully captures features of obstacles that unsupervised learning alone struggles to discern. For instance, roadside barriers(Fig.13, Col 2), which run parallel to the direction of vehicle movement and remain unchanged across multiple frames, pose challenges for unsupervised methods but are accurately predicted with the aid of synthetic depths. Synthetic depths are dense and accurate, providing crucial information about these obstacles.
Furthermore, the characteristics of synthetic datasets are also learned by the network. For example, as shown in Fig.11, the SHIFT dataset has two characteristics compared to the vKitti2. Firstly, it contains a amount of pedestrians, which are often difficult to estimate correctly in unsupervised depth estimation, typically resulting in partial depth disappearance or blurring, as seen in the estimation results of DIFFNet in the Fig.15, Row 1. When using our method with the SHIFT dataset, the network learns the knowledge of estimating pedestrian, achieving clear and accurate estimation of pedestrian depth, which is impressive.
This point can also be inferred from predicted depth of transparent objects, such as glass. As shown in Fig.11, in the vKitti2, transparent objects like glass are treated as solid entities and can be detected, resulting in smooth surfaces in Depth. However, in the SHIFT, transparent objects are considered penetrable targets, leading to discontinuous depth in these regions. This difference in Synthetic dataset definitions is perfectly reflected in the training, as shown in Fig.16, Row 2. Models trained with the SHIFT dataset recognize and penetrate through vehicle windows, whereas those trained with the vKitti2 dataset perceive them as flat surfaces.
Figure 17: An experiment predicting color from ground truth (GT) depth showed that shadows and road markings vanished, undetectable in the depth image. See Discussion for details.
IV-F Discussion
IV-F1 Why can the network transform information beyond depth
in Fig. 3, Syn-to-Real images exhibit excessive detail surpassing information in depth, such as tree shadows and lane markings. We attribute this to the Color-Depth-Color process, where depth, predicted from color, retains texture details (such as subtle gradient changes at some boundaries). Consequently, the Color Net can recover such gradient changes from the depth map. To validate, we directly trained the network to predict color images from synthesized depth in the synthetic dataset (Fig. 17). The absence of lane markings and shadows in the restored color images confirms our hypothesis.
IV-F2 Why learn Depth and Color Net simultaneously online
the primary objective of our method is depth prediction. As the Depth Net undergoes continuous updates to enhance its accuracy and performance, it is imperative to ensure that the Color Net also adapts accordingly. This synchronization is crucial for achieving superior color transformation effects. By reprojecting the Color Net’s adjustments with the updates in the Depth Net, we aim to optimize the overall performance of the model. This dynamic adjustment process ensures that the color transformation remains effective and reproject with the evolving depth predictions.
IV-F3 Can the VAN-based encoder replace the ResNet structure in the Pose encoder
so far, we have not observed any new research that can replace the ResNet-based PoseEncoder[9], whether using CNN-based or Transformer-based methods. We have attempted some alternatives, but none have improved depth estimation performance; in fact, many have resulted in performance degradation. In this study, we also tried using a VAN-based model to replace the ResNet-based PoseEncoder, but still did not observe any significant improvement. We believe that the pose estimation task is a small parameter regression task that is relatively simple compared to depth prediction and does not require a powerful network.
V Conclusion
In this paper, we presented Back to the Color(Back2Color), a novel framework designed to enhance unsupervised depth estimation by leveraging synthetic data. Our method addresses the challenge of aligning synthetic world with real world colors, enabling the transformation of predicted synthetic depth into realistic colors with high fidelity. By joint training on synthetic and real-world datasets using Back2Color, we observed substantial improvements in depth estimation performance. Additionally, we designed a depth net based on the Visual Attention Network (VAN), which allowed our CNN-based network to surpass several Transformer-based depth models. Furthermore, we improved the uncertainty-based loss fusion method by combining the advantages of temporal and spatial reprojection, thoroughly addressing issues related to non-rigid motion. Experimental results demonstrated the effectiveness of our method on both the KITTI and Cityscapes datasets. Overall, Back2Color offers a promising solution for advancing unsupervised depth estimation by effectively integrating synthetic data into the training process.

### v7

@@
-Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation
+Back2Color: Domain-Adaptive Synthetic-to-Real Monocular Depth Estimation for Dynamic Traffic Scenes
 Abstract
-Virtual engines can generate dense depth maps for various synthetic scenes, making them invaluable for training depth estimation models.
-However, discrepancies between synthetic and real-world colors pose significant challenges for depth estimation in real-world scenes, especially in complex and uncertain environments encountered in unsupervised monocular depth estimation tasks.
-To address this issue, we propose Back2Color, a framework that predicts realistic colors from depth using a model trained on real-world data, thus transforming synthetic colors into their real-world counterparts.
-Additionally, we introduce the Syn-Real CutMix method for joint training with both real-world unsupervised and synthetic supervised depth samples, enhancing monocular depth estimation performance in real-world scenes.
-Furthermore, to mitigate the impact of non-rigid motions on depth estimation, we present an auto-learning uncertainty temporal-spatial fusion method (Auto-UTSF), which leverages the strengths of unsupervised learning in both temporal and spatial dimensions.
-We also designed VADepth, based on the Visual Attention Network, which offers lower computational complexity and higher accuracy than transformers.
-Our Back2Color framework achieves state-of-the-art performance on the Kitti dataset, as evidenced by improvements in performance metrics and the production of fine-grained details.
-This is particularly evident on more challenging datasets such as Cityscapes for unsupervised depth estimation.
-Code is available at https://github.com/BlueEg/back2color/.
-Index Terms:
-Back2Color, Syn-Real CutMix, auto-learning uncertainty, Visual Attention Depth Network
+Accurate monocular depth estimation is a fundamental component of vision-based perception systems in intelligent transportation applications.
+Despite recent progress, unsupervised monocular approaches still suffer from significant performance degradation in real-world traffic scenes due to synthetic-to-real domain gaps and the presence of dynamic, non-rigid objects such as vehicles and pedestrians.
+In this paper, we propose Back2Color, a robust unsupervised monocular depth estimation framework that addresses these challenges through domain adaptation and uncertainty-aware fusion.
+Specifically, Back2Color proposes a bidirectional depth-to-color transformation strategy that learns appearance mappings from real-world driving data and applies them to synthetic depth maps, thereby constructing training samples with realistic color appearance and paired synthetic depth.
+In this way, the proposed approach effectively reduces the domain gap between simulated and real traffic scenes, enabling the depth prediction network to learn more stable and generalizable priors.
+To further improve robustness under dynamic environments, we propose an auto-learning uncertainty temporal-spatial fusion (Auto-UTSF) module, which adaptively fuses complementary temporal and spatial cues by estimating pixel-wise uncertainty, enabling reliable depth prediction in the presence of moving objects and occlusions.
+Extensive experiments on challenging urban driving benchmarks, including KITTI and Cityscapes, demonstrate that the proposed method consistently outperforms existing unsupervised monocular depth estimation approaches, particularly in dynamic traffic scenarios, while maintaining high computational efficiency.
 Figure 1: Performance vs.
@@
 In the real world, even after prolonged evolution, humans are still limited to qualitatively analyzing the depth of specific points in their field of view to assist in their actions[23].
-Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[4, 5].
+Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[15, 36, 30, 50, 39, 37].
 Even with the refinement of advanced LiDAR systems[36], their data are inherently sparse, comprising only discrete depth points.
-Unsupervised depth estimation(UDE), powered by epipolar geometry[6] using purely visual approaches does not require additional annotation for depth information other than camera parameters.
+Unsupervised depth estimation(UDE), powered by triangulation[42] using purely visual approaches does not require additional annotation for depth information other than camera parameters.
 UDE eliminates the necessity of annotating depth labels for specific scenes.
@@
 Additionally, sparse depth sampling does not provide sufficient information for learning when obstacles are small or distant (resulting in small apparent areas), making it challenging for the network to accurately reconstruct their shapes.
-Figure 2: Illustration of our framework.
-Our Back2Color training follows a two-step optimization process within each iteration.
-In the first step, we use Optimizer 1 to train the Depth Net for supervised Syn-Real depth estimation and the Color Net to predict real colors from depths.
-Initially, ① we freeze the Depth Net and Color Net to infer both real and synthetic predicted depths, using synthetic depth data to predict Syn-to-Real Color.
-Next, ② we apply Syn-Real CutMix to generate new syn-real colors, and ③ optimize the Depth Net with synthetic ground truth and the Color Net on the real dataset.
-In the second step, with Optimizer 2, ④ we train for real-world unsupervised depth estimation, incorporating sky regularization loss and auto-learning uncertainty techniques.
-Note that the dotted line indicates inference only.
-In this paper, our Back2Color learns the mapping process in the target domain (Real).
-This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains.
-Additionally, we introduced an improved version of CutMix[20], termed Syn-Real CutMix, to effectively combine real-world, synthesized, and syn-to-real data in both batch and image spatial dimensions.
-By incorporating Syn-Real CutMix into our training process, we were able to enrich the training dataset with more diverse color-to-depth knowledge extracted from synthetic scenes.
-Furthermore, the joint training with diverse data sources enables the model to capture a wider range of color and depth variations present in real-world scenes, thus enhancing its robustness and generalization capability.
-Additionally, aiming to narrow the gap between predicted depths from real-world and synthesized depth, we addressed the issue of chaotic depth clouds (5) by introducing a Background Textureless Regularization Loss, also known as Sky Regularization Loss.
-Through joint training, our method enables the unsupervised monocular depth estimation to leverage knowledge from synthesized data, resulting in the generation of dense and sharp depths.
-In pursuit of higher prediction accuracy and a more lightweight model compared to Transformer-based methods, we introduced the Visual Attention Network[21](VAN) as the backbone and equipped it with an efficient single-path high-resolution decoder, forming the new Visual Attention Depth Net(VADepth).
-Our design demonstrated more effective performance than Transformers.
-In our in-depth experiments, we found that Auto-masking is not always effective in identifying non-rigid moving objects with the same direction and speed.
-Additionally, it cannot handle vehicles moving in opposite directions, leading to significant errors in depth estimation.
-Therefore, we introduce the automatic learning uncertainty estimation fusion method[22], effectively combining the advantages of temporal-spatial unsupervised approaches, termed Auto-learning Uncertainty Temporal-Spatial Fusion(Auto-UTSF).
-This method thoroughly resolves the issue of non-rigid motion in unsupervised estimation, greatly improving the accuracy of depth estimation.
-To summarize, the key contributions of our work include:
-  • To address the issue of color and texture discrepancies between synthetic and real-world datasets, we propose Back2Color.
-This framework stores the color information of specific scenes (target domain) within the model, enabling the reverse prediction of color images from depth data.
-Additionally, we introduce Syn-Real CutMix to harmonize the joint training of real-world and synthetic data, thereby enhancing the effectiveness of unsupervised depth estimation in real-world scenes.
-  • We designed an efficient and lightweight Depth Net based on the Visual Attention Network.
-This network boasts fewer parameters and lower computational complexity compared to transformer-based networks, while delivering superior performance.
-  • To more robustly recognize non-rigid moving objects and address occlusion issues in stereo systems, we enhance the method based on Auto-Uncertainty by introducing Auto-Spatial-Temporal Uncertainty Fusion, which leverages complementary spatial and temporal advantages in the unsupervised depth estimation process.
-  • Experimental results demonstrate that our method surpasses all transformer-based models, achieving state-of-the-art (SOTA) metrics while maintaining manageable computational costs.
-Additionally, the visual boundaries of the estimated depth images are clearer.
-Figure 3: Color transformation for specific datasets: we showcase the capability of our Back2Color framework to effectively transform various synthetic datasets, including vKitti2[12], SHIFT[23], and UrbanSyn[24], into Syn2Real images.
+Figure 2: Illustration of the Back2Color training strategy.
+In each iteration, the first stage jointly optimizes the depth estimation network and the depth-to-color network using real-world data with unsupervised photometric and color restoration losses.
+In the second stage, the depth-to-color network is fixed and applied to transform synthetic depth maps into realistic color images, enabling the depth estimation network to be trained with virtual depth supervision under a reduced domain discrepancy.
+In this paper, we propose the Back2Color framework to address the appearance discrepancy between synthetic and real-world data in unsupervised monocular depth estimation.
+Specifically, Back2Color introduces a bidirectional depth-to-color transformation strategy that learns appearance mappings from real-world driving scenes and applies them to synthetic depth maps.
+In this way, Back2Color constructs training pairs that combine realistic color appearance with corresponding synthetic depth, effectively narrowing the domain gap between simulated and real traffic scenes.
+To further enable effective joint training across heterogeneous data sources, we introduce an enhanced CutMix strategy, termed Syn-Real CutMix [59], which mixes synthetic and syn-to-real images at both batch and spatial levels.
+This strategy enriches the training data with diverse color–depth associations derived from synthetic scenes while preserving real-world appearance characteristics.
+By jointly training on these complementary data sources, the proposed framework allows the depth estimation network to learn more stable and generalizable depth prediction priors, leading to improved robustness in real traffic environments.
+In addition, we observe that existing auto-masking mechanisms often fail to reliably identify non-rigid moving objects in traffic scenes, especially when vehicles move at similar speeds or in opposite directions, which leads to erroneous depth predictions.
+To address this issue, we introduce an automatic uncertainty learning fusion strategy that combines temporal and spatial cues, termed Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) [69].
+By explicitly modeling uncertainty, Auto-UTSF adaptively integrates complementary information from stereo and temporal supervision, enabling more reliable depth estimation under dynamic and occluded conditions.
+To summarize, the main contributions of this work are as follows:
+  • We propose Back2Color, a synthetic-to-real adaptation framework that learns scene-specific color information from real-world data and enables reverse color prediction from depth.
+Together with the proposed Syn-Real CutMix strategy, Back2Color effectively harmonizes joint training on synthetic and real datasets, improving the robustness of unsupervised monocular depth estimation in real traffic scenes.
+  • We introduce Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF), which robustly handles non-rigid motion and occlusions by adaptively fusing complementary spatial and temporal supervision through uncertainty estimation.
+  • Extensive experiments on challenging urban driving benchmarks demonstrate that the proposed method consistently improves depth estimation accuracy while maintaining computational efficiency, and produces depth maps with clearer object boundaries.
+Figure 3: Color transformation for specific datasets: we showcase the capability of our Back2Color framework to effectively transform various synthetic datasets, including vKITTI2[4], SHIFT[51], and UrbanSyn[gómez2023one], into Syn2Real images.
 These images closely emulate the real-world styles of KITTI[10, 12] and Cityscapes[7], demonstrating high fidelity and credibility.
@@
 Initial methods like those by Godard et al.[13] introduced disparity consistency loss, while Zhou et al.[67] pioneered depth and pose estimation simultaneously.
-Recent advancements[28, 29, 30, 31, 32, 33, 34, 35] refine network structures for higher accuracy.
+Recent advancements[18, 57, 43, 66, 22, 38] refine network structures for higher accuracy.
 Techniques like PackNet[18] and CADepth[57] enhance geometric and appearance information propagation.
@@
 Zheng et al.[65] used GANs to convert synthetic colors to real-world colors.
+With advancements in stable diffusion[61, 53], ControlNet [61] is now capable of generating images based on various conditions, including text, pose, Canny edges, and depth.
+This capability theoretically enables the generation of RGB images from urban depth maps, as demonstrated by RISHI et al.
+[53].
+However, experiments [53] indicate that diffusion models do not strictly adhere to depth information and may arbitrarily alter target classes.
+Additionally, diffusion models require 2-5 seconds to generate a single image espically trainning with CNN model, and each condition demands extensive data for training.
+This makes it challenging to quickly fine-tune a depth estimation model for specific real-world scenarios.
+In contrast, our proposed ColorNet addresses these limitations, allowing for rapid deployment and real-time operation.
 III Methodology
-In this section, we introduce the Back2Color framework, which masters the depth-to-color mapping within the target domain to train the Color Net (III-A1).
-Back2Color is enhanced by Real-Syn CutMix (III-A2) and a strategy for handling textureless background objects (III-A3).
-We then discuss the VADepth, based on the Visual Attention Network, designed for high performance and lightweight architecture (III-B).
-Lastly, we describe the Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) (III-C) for addressing non-rigid motion challenges.
-In this paper, Depth and Disparity are equivalent.
+In this section, we present the Back2Color framework, which captures the depth-to-color mapping within the target domain and transfers synthetic colors to pseudo-realistic ones (III-A1).
+Back2Color is further enhanced with Real-Syn CutMix (III-A2) to jointly train for unsupervised depth estimation in the real world and supervised depth estimation in synthetic data.
+Then, we describe the Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) module (III-B), designed to address challenges associated with non-rigid motion.
+Throughout this paper, Depth and Disparity are considered equivalent.
 Figure 4: Syn-Real CutMix: specific steps involved in our Syn-Real CutMix method for spatially combining colors: (a) Randomly select a rectangular region, denoted as M.
@@
 The colors in real word datasets like Cityscapes and KITTI are notably different, as shown in Fig.3.
-Therefore, in order to bridge the substantial gap between synthetic $\mathcal{S}\in${vKitti2[12](Vk),SHIFT[23](S), UrbanSyn[24]} and real-world $\mathcal{X}\in$Kitti[25, 19](K),Cityscapes[11](Cs)}datasets, we can learn the depth-to-color mapping in the target domain (real world color) and apply it to the source domain (synthetic color), thereby converting the source domain color to match those in the target domain.
-And then, we made the following attempts, train a Color Net $\mathcal{F}_{d2c}$ to predict color images $\tilde{\mathbf{I}}^{\mathcal{X}}$ from predicted depth images $\hat{\mathbf{D}}^{\mathcal{X}}$:
-$\hat{\textbf{I}}^{\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\textbf{D}}^{\mathcal{X}}|\boldsymbol{\theta}^{\mathcal{X}}),$ (1)
-and utilizing the input of the Depth Net $\textbf{I}^{\mathcal{X}}$ for Supervision:
-$\mathcal{L}^{d2c}=pe(\hat{\textbf{I}}^{\mathcal{X}},\textbf{I}^{\mathcal{X}}),$ (2)
-Figure 5: Chaotic depth clouds: The green pixels can be matched across relevant frames based on the features of foreground obstacle, while the red pixels lacks effective features for matching, resulting in a cloud-like region in the depth image.
-where $\boldsymbol{\theta}^{\mathcal{X}}$ represents the weights that store specific dataset color mapping information, such as Kitti and Cityscapes.
-Fortunately, we found that this prediction is feasible, neural networks are able to capture color information under specific target scenes.
-Next, we need to investigate whether we can employ $\boldsymbol{\theta}^{\mathcal{X}}$ to convert predicted $\hat{\textbf{D}}^{\mathcal{S}}$ in synthesized datasets into color that resemble the Kitti($\mathcal{X}$) color style: Using the specific Color Net to colorize predicted synthetic depth, as shown in Fig.3:
-$\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\textbf{D}}^{\mathcal{S}}|\boldsymbol{\theta}^{\mathcal{X}}).$ (3)
-Figure 6: The proposed depth estimation module based on the Visual Attention Network.
-The encoder is composed of an ImageNet pretrained Visual Attention Network (VAN), generating multi-scale features.
-The VAN stages consist of the Large Kernel Attention (LKA) module, which is a pure CNN structure.
-This encoder balances performance and efficiency.
-The decoder features an HRNet structure, where each stage fuses multi-scales of features upward.
-Our experiments have shown that by referencing color information from the target domain, we can learn its depth-to-color mapping and accurately predict color images in the source domain that closely resemble those in the target domain.
+Therefore, in order to bridge the substantial gap between synthetic $\mathcal{S}\in$ {vKITTI2(vK)[4], SHIFT(SH)[51], UrbanSyn(Ur)[gómez2023one]} and real-world $\mathcal{X}\in$ {KITTI(KT)[10, 12], Cityscapes(Cs)[7]} datasets, we can learn the depth-to-color mapping in the target domain (real world color) and apply it to the source domain (synthetic color), thereby converting the source domain color to match those in the target domain.
+To bridge the appearance gap between synthetic and real-world data, we train a depth-to-color network $\mathcal{F}_{d2c}$ that predicts color images from estimated depth maps in the real-world domain.
+Given a predicted depth map $\hat{\mathbf{D}}_{\mathcal{X}}$ produced by the depth estimation network, the depth-to-color network reconstructs the corresponding color image as
+$\hat{\mathbf{I}}_{\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\mathbf{D}}_{\mathcal{X}};\boldsymbol{\theta}_{d2c}),$ (1)
+where $\boldsymbol{\theta}_{d2c}$ denotes the learnable parameters of the depth-to-color network.
+The reconstructed color image $\hat{\mathbf{I}}_{\mathcal{X}}$ is supervised by the corresponding real color image $\mathbf{I}_{\mathcal{X}}$ using a photometric reconstruction loss, defined as
+$\mathcal{L}_{d2c}=pe(\hat{\mathbf{I}}_{\mathcal{X}},\mathbf{I}_{\mathcal{X}}),$ (2)
+where $pe(\cdot,\cdot)$ denotes the photometric error function.
+Next, we investigate whether the depth-to-color mapping learned from the real-world domain can be transferred to synthetic data.
+Specifically, after training the depth-to-color network $\mathcal{F}_{d2c}$ on real-world samples, we apply the learned model to predicted depth maps from the synthetic domain in order to generate color images that resemble the appearance characteristics of the target domain (e.g., KITTI), as illustrated in Fig.
+3.
+Formally, given a predicted synthetic depth map $\hat{\mathbf{D}}_{\mathcal{S}}$, the corresponding color image is generated as
+$\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}}=\mathcal{F}_{d2c}(\hat{\mathbf{D}}_{\mathcal{S}};\boldsymbol{\theta}_{d2c}),$ (3)
+where $\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}}$ denotes the synthetic color image transformed to match the appearance of the real-world domain.
 III-A2 Syn-Real CutMix
-CutMix[20] replaces removed pixels with patches from other images rather than simply removing them.
-This strategy eliminates uninformative pixels during training, enhancing efficiency, while maintaining the advantages of regional dropout to focus on non-discriminative obstacle parts.
-Additionally, introducing extra patches improves the model’s localization ability, requiring identification of obstacles from partial views.
-In our work, we draw inspiration from the CutMix method to achieve different objectives and effects.
-We aim to dynamically fuse Syn2Real Colors and Synthetic Colors within a single batch for supervised training, a technique we have named Syn-Real CutMix.
-The goal of CutMix in our work is to generate a new training pair $(\tilde{\textbf{I}},\textbf{D})$.
-Unlike CutMix[20], our samples$(\textbf{I}^{\mathcal{S}},\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}})$ are not two unrelated samples; instead, they correspond to different colored images of the same depth label $\textbf{D}^{\mathcal{S}}$.
-Therefore, we only merge color images without combining new labels.
-To increase the randomness of training samples.
-As shown in Fig.4, given image samples $(\textbf{I}^{\mathcal{S}},\hat{\textbf{I}}^{\mathcal{S}\rightarrow\mathcal{X}})$ as $(\textbf{I}_{1},\textbf{I}_{2})$, we firstly randomly swap $(\textbf{I}_{1},\textbf{I}_{2})$.
-Then, randomly crop a block from $\textbf{I}_{1}$ with a height and width greater than half of the original image, and replace it with the corresponding position in $\textbf{I}_{2}$:
-$\tilde{\textbf{I}}=\textbf{M}\odot\textbf{I}_{1}+(1-\textbf{M})\odot\textbf{I}_{2},$ (4)
-where $M_{i}\in\{0,1\}$ denotes a binary mask indicating where to drop out and fill in, and $\odot$ is element-wise multiplication.
-Now, we combine Syn-Real enhanced images $\tilde{\textbf{I}}$ and Real world target image I into one batch for training.
-Then, we can train monocular depth estimation networks simultaneously on real-world, synthetic, and Syn-to-Real images within a batch.
-Spatial randomization also encourages the network to focus more on the image structure rather than its specific colors.
-III-A3 Background textureless regularization loss
-when analyzing and comparing unsupervised monocular depth and synthesized depth, we found not only significant gap in color images but also substantial gaps in the depth outputs generated by the model.
-In Fig.5, when the background of obstacle is textureless regions, these areas lack features for matching in sequences or stereo, making it challenging to accurately determine their depth.
-Consequently, these regions are prone to arbitrary depth assignments within the textureless area, leading to the formation of cloud-like regions in the depth map, we called them chaotic depth clouds.
-The main drawback of such cloud-like regions lies in their ability to significantly impact the overall distribution of depth values in the depth map, resulting in notable biases when calculating median variance during evaluation processes.
-For chaotic depth clouds, they appear mostly within regions with the sky as background, such as trees, road signs, and buildings.
-Fortunately, the sky’s depth itself has a strong prior, that it is infinitely far away.
-Therefore, we predict the mask of the sky and constrain its value on the disparity map estimation by the network.
-This ensures that the network separates background and foreground obstacles:
-$\mathcal{L}^{sky}=\frac{\sum|\textbf{M}_{sky}\times\hat{\textbf{D}}|}{\sum\textbf{M}_{sky}},$ (5)
-where $\textbf{M}_{sky}$ is inferred from a pre-trained semantic segmentation network[52].
-We also have an intuitive prior that the sky should only exist above in the view, so we added a decay coefficient $\textbf{G}\in\mathbb{R}^{W\times H}$ that decreases with the image height:
-$\textbf{G}[..,j]=(H-j)/H,$ (6)
-where $H$ is the height of $\textbf{D}_{t}$, so our final sky regularization loss function is:
-$\mathcal{L}^{sky}=\frac{\sum\textbf{G}\times|\textbf{M}_{sky}\times\hat{\textbf{D}}|}{\sum\textbf{M}_{sky}}$ (7)
-The proposed sky regularization loss effectively eliminates hazy depth while enhancing the recovery of delicate objects.
+The objective of CutMix in our framework is to construct enhanced training pairs $(\tilde{\mathbf{I}},\mathbf{D})$ for improving robustness across domains.
+Unlike the original CutMix [59], which mixes two independent image–label pairs, the image pair used in our setting, $(\mathbf{I}_{\mathcal{S}},\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}})$, represents two different color appearances corresponding to the same synthetic depth label $\mathbf{D}_{\mathcal{S}}$.
+Therefore, we perform mixing only in the color space while keeping the depth supervision unchanged.
+To introduce spatial diversity, as illustrated in Fig.
+4, we denote the image pair $(\mathbf{I}_{\mathcal{S}},\hat{\mathbf{I}}_{\mathcal{S}\rightarrow\mathcal{X}})$ as $(\mathbf{I}_{1},\mathbf{I}_{2})$, which are randomly swapped with equal probability.
+A rectangular region with height and width larger than half of the original image is then randomly cropped from $\mathbf{I}_{1}$ and replaced with the corresponding region from $\mathbf{I}_{2}$, yielding the mixed image
+$\tilde{\mathbf{I}}=\mathbf{M}\odot\mathbf{I}_{1}+(1-\mathbf{M})\odot\mathbf{I}_{2},$ (4)
+where $\mathbf{M}\in\{0,1\}^{H\times W}$ denotes a binary mask indicating the preserved and replaced regions, and $\odot$ represents element-wise multiplication.
+The Syn-Real enhanced images $\tilde{\mathbf{I}}$ are then combined with real-world target images $\mathbf{I}_{\mathcal{X}}$ within the same training batch.
+This joint batching strategy allows the monocular depth estimation network to be trained simultaneously on real-world, synthetic, and Syn-to-Real samples.
+The spatial randomization introduced by Syn-Real CutMix further encourages the network to focus on structural cues rather than specific color appearances, thereby improving generalization across domains.
 Figure 5: When network capabilities become stronger, and the number of non-rigid objects in the dataset increases, the shortcomings[69] of Auto-masking and Minimum Reprojection Error lead to the reappearance of errors.
-III-B Visual Attention Depth Net
-Every high-performing depth estimation network requires an effective and robust backbone to extract hierarchical features from natural images.
-Commonly used backbones include ResNet[53] (Monodepth2[9], HRDepth[30], CADepth[29]), HRNet[54] (DIFFNet[31], RA-Depth[32]), and Transformer-based backbones like ViT[37], MPViT[41](TSUDepth[22], MonoViT[2]).
-With the evolution from ResNet to HRNet and then to Transformer-based approaches, the performance of depth estimation continues to improve, but accompanied by increasing computational complexity and params.
-We consistently believe that depth estimation tasks, especially those based on street scenes, need to consider the requirements of their usage scenarios (real-time, lightweight).
-Hence, we aim to find a more efficient and effective backbone than Transformer methods.
-Research on balancing the performance and efficiency of CNN and Transformer models has been ongoing, and we have discovered the VAN[21] (Visual-Attention-Network) model.
-The VAN[21] model employs large kernel convolutions instead of self-attention mechanisms to capture long-range dependencies.
-As shown in Fig.6, It decomposes the large kernel convolution into a spatial local convolution (depth-wise convolution), a spatial long-range convolution (depth-wise dilation convolution), and a channel convolution (1×1 convolution) to reduce computational complexity.
-This design not only avoids treating 2D images as 1D sequences but also achieves stronger accuracy on ImageNet while reducing computational complexity and params.
-Due to its outstanding performance and efficiency, we have introduced the VAN model into unsupervised monocular depth estimation tasks, and adapt to some fine-tuning.
-The most significant change is that, to reduce computational complexity, VAN downsamples the input resolution to $2^{-2}$ during the first stem stage.
-While this downsampling is acceptable for classification tasks, it poses challenges for high-resolution recovery in image regression tasks.
-Therefore, without altering the original VAN structure, we have added an additional stem head to acquire features at $2^{-1}$ resolution, as shown in Fig.6.
-After extracting features from various VAN stages, we equipped them with a high-resolution depth decoder[32] to output the final depth.
-During implementation, we observed no difference between outputting a single-size depth image and outputting multi-scale depth images.
-Therefore, our network only outputs depth maps with the same size as the input.
-Figure 8: Temporal reprojection involves reprojecting the previous and subsequent frames to the target frame, while spatial reprojection involves reprojecting the left or right frame to the target frame.
-III-C Auto-learning uncertainty temporal-spatial fusion
-When testing our designed VADepth model, we found that the Auto-masking[9] method almost failed, as shown in Fig.7.
-This issue is more severe in the Cityscapes[11].
-Upon inspecting the output of Auto-masking, we discovered that it couldn’t perfectly shield non-rigid objects.
-We attribute this to the model’s robustness, the limitations of the Auto-masking method, and the complexity of the Cityscapes[11, 55].
-Therefore, we need to address the issue of non-rigid objects to further enhance performance.
-In temporal sequential frames, there are freely moving targets, such as pedestrians and vehicles(as show in Fig.8).
-These do not conform to the rigid scene transformation assumptions based on depth and camera pose.
-The solution provided by Auto-Masking[9] is to compare the pixels before and after reprojection.
-If the reprojection error after reprojection is worse than without reprojection, it indicates that the pixel belongs to a non-rigid object.
-$\mu=[\min_{c}pe(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})<\min_{c}pe(\textbf{I}_{t},\textbf{I}_{c})].$ (8)
-However, the network can still rationalize the depth of non-rigid moving objects by pulling forward targets further away and pushing backward closer targets, making them consistent with the camera extrinsics(Pose), as show in Fig.7.
-In fact, non-rigid moving objects only exist in temporally related frames, while in spatially related frames (stereo settings), since they are captured at the same moment, this issue does not arise.
-Therefore, introducing stereo supervision into the system is the most direct and effective method, rather than merely excluding masks for non-rigid objects[55, 9].
-Figure 9: Auto-Spatio-Temporal Uncertainty Fusion: The Uncertainty Generation Network utilizes a Pose Encoder and establishes an Uncertainty Decoder.
-It inputs the $\textbf{I}_{\{-1,+1,0,s\}}$ and the target frame $\textbf{I}_{0}$ separately to output the uncertainty map $\hat{\textbf{U}}$.
-Then, it generates the weight map U for different error maps through a “1-Softmax” operation.
-But when comparing temporal sequence reprojection with spatial reprojection, each has its own advantages and disadvantages.
-First, the translation of pose between temporally related frames are larger, whereas in spatial reprojection (stereo), the baseline is fixed and small.
-This makes temporal-based transformations better suited for estimating distant targets.
-Then, temporally related frames typically include three frames: the current frame and the preceding and following frames, allowing information from the preceding and following frames to fill in occluded areas in the target frame.
-In spatial reprojection (stereo), regardless of whether the target frame is set as the left or right eye, there is no symmetric completion of occluded areas.
-Thus, we seek a method that can effectively combine the advantages of both temporal and spatial reprojections to predict better depth maps.
-We referenced prior studies on Uncertainty Learning[56, 57, 22].
-And in the work by Zhu et al.[22], a method was introduced to address occlusion problems by automatically predicting uncertainty and reweighting the reprojection in the reprojection process of spatially adjacent frames.
-Building on this idea, we aim to take it a step further by introducing a method to predict uncertainty during both temporal and spatial reprojections.
-This approach seeks to leverage the strengths of both temporal and spatial domains effectively.
-To combine photometric loss across different frames, Zhu et al.[22] specify $\sum_{c}(1-\hat{u}_{i,c})=1,c\in\{t-1,t+1\}$.
-Hence, they combine the output uncertainty map $\hat{\textbf{U}}_{c}(\hat{\textbf{U}}_{c}\geq 0)$ with softmax: $\{\textbf{U}_{t-1},\textbf{U}_{t+1}\}=1-Softmax\{\hat{\textbf{U}}_{t-1},\hat{\textbf{U}}_{t+1}\}$, to ensure that $\sum_{c}u_{i,c}=1$.
-The final joint optimization function is written as:
-$\begin{split}\mathcal{L}&=\frac{1}{N}\sum^{N}\sum_{c}u_{i,c}pe_{i}(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})\\ &\quad+\alpha\frac{1}{N}\sum^{N}\sum_{c}\hat{u}_{i,c},\quad\sum_{c}u_{i,c}=1.\end{split}$ (9)
-Using this formula, the network will automatically learn the uncertain regions(areas where the loss cannot be reduced through optimization) within the loss, forming the final weights for multi-frame loss through comparison.
-In the temporal multi-frame reprojection loss combination, high-uncertainty regions are the frames’ occluded areas revealed due to reprojection.
-Meanwhile, in the spatial and temporal reprojection loss combination, high-uncertainty regions in the temporal reprojection loss map are the non-rigid motion areas that can be accurately reprojected in the spatial.
-$\begin{split}\mathcal{L}^{uD}=\frac{1}{N}\sum^{N}\Bigg{(}u_{i,\tau}\sum_{c}u_{i,c}pe_{i}(\textbf{I}_{t},\textbf{I}_{c\rightarrow t})+u_{i,s}pe_{i}(\textbf{I}_{t},\textbf{I}_{s\rightarrow t})\Bigg{)}\\ +\alpha\frac{1}{N}\sum^{N}(\sum_{c}\hat{u}_{i,c}+\sum_{\sigma}\hat{u}_{i,\sigma}),\quad c\in\{-1,+1\},\quad\sigma\in\{\tau,s\}\end{split}$ (10)
-where $\tau$ and $s$ represent the temporal domain and the spatial domain, respectively, with $\alpha$ set to $1e^{-2}$.
-$u_{i,\tau}$ and $u_{i,s}$ represent the temporal-domain uncertainty and the spatial-domain uncertainty, $\sum_{c}u_{i,c}=1,\sum_{\sigma}u_{i,\sigma}=1$, $\{\textbf{U}_{\tau},\textbf{U}_{s}\}=1-Softmax\{\hat{\textbf{U}}_{\tau},\hat{\textbf{U}}_{s}\}$.
-$s$ represents the source frame in a stereo setup.
-Since the Kitti dataset randomly selects either the left or right camera image as the target frame, the other frame becomes the source frame.
+Figure 6: Illustrates the complementary characteristics of temporal and spatial reprojection, which motivates the proposed Auto-UTSF strategy.
+III-B Auto-learning Uncertainty Temporal–Spatial Fusion
+During the evaluation of the Back2Color framework, we observe that the widely adopted auto-masking strategy [14] often fails in complex urban scenes, particularly on the Cityscapes dataset, as illustrated in Fig.
+5.
+By inspecting the auto-masking outputs, we find that non-rigid moving objects, such as vehicles and pedestrians, cannot be reliably filtered.
+This limitation is further aggravated by the complexity of real-world traffic scenes and the strong fitting capability of modern depth networks, which may compensate for non-rigid motion by adjusting depth predictions to remain consistent with camera poses [27].
+Moreover, masking strategies based on optical flow [69] completely fail for objects moving in the opposite direction of the ego-motion, as shown in Fig.
+5.
+In temporal sequences, freely moving objects violate the rigid-scene assumption underlying depth- and pose-based reprojection, as depicted in Fig.
+6.
+Auto-masking attempts to identify such pixels by comparing reprojection errors before and after geometric warping:
+$\mu_{i}=\mathbb{I}\!\left[\min_{k}pe(\mathbf{I}_{t},\mathbf{I}_{k\rightarrow t})<\min_{k}pe(\mathbf{I}_{t},\mathbf{I}_{k})\right],$ (5)
+where $k\in\{t-1,t+1\}$. However, this heuristic is insufficient when depth predictions can be adjusted to rationalize non-rigid motion, leading to incorrect supervision.
+Notably, non-rigid motion only affects temporally adjacent frames, whereas in spatially related stereo pairs captured at the same time, such motion does not occur.
+Therefore, instead of simply excluding unreliable pixels, introducing spatial (stereo) supervision provides a more direct and effective solution.
+Nevertheless, temporal and spatial reprojections exhibit complementary characteristics.
+Temporal reprojection benefits from larger baselines and multi-frame context, making it suitable for distant regions and occlusion completion, while spatial reprojection offers reliable supervision for non-rigid objects but suffers from limited baseline and asymmetric occlusions.
+These observations motivate a unified strategy that adaptively fuses temporal and spatial supervision.
+Inspired by uncertainty learning [68, 69, 44], we introduce an Auto-learning Uncertainty Temporal–Spatial Fusion (Auto-UTSF) mechanism to combine temporal and spatial reprojection losses.
+In this section, subscripts $t$, t1t\!-\!1, t+1t\!+\!1, and $s$ are used to index temporal and stereo frames, while dataset/domain subscripts (e.g., $\mathcal{X}$ and $\mathcal{S}$) are omitted for clarity when the context is unambiguous.
+And $i$ indexes individual pixels in the image, and $N$ denotes the total number of pixels.
+Following [69], pixel-wise temporal uncertainty logits $\hat{u}^{\tau}_{i,k}$ are predicted for temporally adjacent frames $k\in\{t-1,t+1\}$ and normalized using a softmax operation:
+$u^{\tau}_{i,k}=1-\text{Softmax}(\hat{u}^{\tau}_{i,t-1},\hat{u}^{\tau}_{i,t+1}),\quad\sum_{k}u^{\tau}_{i,k}=1.$ (6)
+And The temporal reprojection loss is then defined as
+$\mathcal{L}_{\tau}=\frac{1}{N}\sum_{i}\sum_{k}u^{\tau}_{i,k}\,pe(\mathbf{I}_{t},\mathbf{I}_{k\rightarrow t})+\alpha\frac{1}{N}\sum_{i}\sum_{k}\hat{u}^{\tau}_{i,k},$ (7)
+where $\alpha$ is set to $1\times 10^{-2}$.
+To further fuse temporal and spatial supervision, we introduce domain-level uncertainty to balance their contributions.
+Temporal-domain and spatial-domain uncertainty logits $(\hat{u}^{\tau}_{i},\hat{u}^{s}_{i})$ are normalized as
+$u^{\tau}_{i},\;u^{s}_{i}=1-\text{Softmax}(\hat{u}^{\tau}_{i},\hat{u}^{s}_{i}),\quad u^{\tau}_{i}+u^{s}_{i}=1.$ (8)
+The final uncertainty-aware temporal–spatial fusion loss is formulated as
+$\mathcal{L}_{\text{ud}}=\frac{1}{N}\sum_{i}\Big($ $u^{\tau}_{i}\sum_{k}u^{\tau}_{i,k}\,pe(\mathbf{I}_{t},\mathbf{I}_{k\rightarrow t})$ (9)
+$+u^{s}_{i}\,pe(\mathbf{I}_{t},\mathbf{I}_{s\rightarrow t})\Big)$
+$+\alpha\frac{1}{N}\sum_{i}\Big(\sum_{k}\hat{u}^{\tau}_{i,k}+\hat{u}^{\tau}_{i}+\hat{u}^{s}_{i}\Big),$
+where $\mathbf{I}_{s}$ denotes the stereo source image corresponding to the target frame $\mathbf{I}_{t}$.
+Figure 7: Auto-Spatio-Temporal Uncertainty Fusion: Given temporal and stereo frame pairs, the uncertainty generation module predicts pixel-wise uncertainty logits, which are normalized via a $1-\mathrm{Softmax}$ operation to adaptively weight temporal and spatial reprojection errors.
+Through uncertainty learning, the model automatically assigns low weights to unreliable regions in each domain.
+In temporal reprojection, high uncertainty mainly corresponds to occlusions and non-rigid motion, while in spatial reprojection, uncertainty primarily arises from occluded regions caused by viewpoint changes.
+By fusing uncertainties from both temporal and spatial domains, Auto-UTSF selectively suppresses unreliable supervision and exploits complementary geometric cues, enabling robust depth estimation in dynamic traffic scenes.
 Figure 8: Temporal Uncertainty for non-rigid objects: We can identify non-rigid objects by predicting the uncertainty during the training process, thereby eliminating their impact in the trainning.
 The fine-grained recognition can even identify very small non-rigid targets.
-Regarding how to obtain temporal and spatial uncertainty maps, we consider referencing the method used to generate temporal uncertainty maps by inputting related frames containing uncertainty information.
-For spatial reprojection, we can input $\textbf{I}_{s}$ and $\textbf{I}_{0}$, but there is no single related frame that uniquely represents temporal reprojection.
-We attempt to input $\textbf{I}_{0}$ and $\textbf{I}_{0}$ as representatives.
-By comparing the output uncertainty maps when calculating the loss, we can form reasonable uncertainty maps, as shown in Fig.
-9.
-Applying uncertainty learning in temporal reprojection can address occlusion and field contraction issues.
-We improved this method to fuse temporal and spatial reprojection errors.
-Since non-rigid objects are rigid in the spatial domain, the temporal domain generates significant uncertainty(Fig.10).
-Additionally, high uncertainty still arises from occluded regions during the spatial domain reprojection process.
-By fusing the uncertainties from both the temporal and spatial domains, we can ignore high-uncertainty areas in each domain when calculating the loss, leveraging the strengths of each domain to train better depths.
-III-D Training Strategy
-III-D1 Synthetic supervised depth loss
+III-C Training Strategy
+III-C1 Synthetic supervised depth loss
 Since unsupervised depth estimation models based on video sequences cannot recover scale information, we adopt the affine-invariant mean absolute error loss used by MiDas[47] and Depth Anything[58] for joint training with synthesized depth, which has deterministic values varying with different camera parameters:
-$\mathcal{L}^{sD}=\frac{1}{HW}\sum_{i}\ |d^{*}_{i}-\hat{d}^{*}_{i}|,$ (11)
+$\mathcal{L}_{sd}=\frac{1}{HW}\sum_{i}\ |d^{*}_{i}-\hat{d}^{*}_{i}|,$ (10)
 where $d^{*}_{i}$ and $\hat{d^{*}_{i}}$ are the scaled and shifted pixels of predicted $\hat{\textbf{D}}$ and synthetic goundtruth D.
@@
 While the MixSyn dataset is trained once, the real-world unsupervised datasets may be trained multiple times.
-III-D2 Optimization Strategy
-Swami et al.[43] highlighted that multiple networks and optimization losses in unsupervised domain adaptation methods often compete against each other and fail to contribute effectively to the optimization of the depth estimation.
-This lack of cooperation among different components makes it challenging to achieve best results.
-So here, we have set two optimizers $\{\mathcal{O}_{d2c},\mathcal{O}_{d2c}\}$ to separately train the Depth Net and the Color net, as shown in Fig.2.
-Optimizer 1-$\mathcal{O}_{c2d}$: Supervised training of the Depth Net on the Syn-Real samples and the Color Net on the Real world unsupervised depth dataset for predicting color from depth, $\gamma=1e^{-3}$.
-$\mathcal{L}^{\mathcal{O}_{d2c}}=\mathcal{L}^{d2c}+\mathcal{L}^{sD}+\gamma\mathcal{L}^{sm}.$ (13)
-Optimizer 2-$\mathcal{O}_{c2d}$: Unsupervised training of the Depth Net on the Real world and applyed the background textureless regularization loss , $\beta=1e^{-2},\gamma=1e^{-3}$.
-$\mathcal{L}^{\mathcal{O}_{c2d}}=\mathcal{L}^{uD}+\beta\mathcal{L}^{sm}+\gamma\mathcal{L}^{sky}.$ (14)
+III-C2 Optimization Strategy
+Swami et al.
+[52] pointed out that, in unsupervised domain adaptation methods, multiple networks and optimization objectives often compete with each other, thereby failing to jointly contribute to effective depth estimation.
+Such a lack of cooperation among different components makes optimization challenging and limits overall performance.
+To address this issue, we adopt two separate optimizers to decouple the training process and independently optimize the Depth Network on synthetic data and real-world unsupervised data, respectively, as illustrated in Fig.
+2.
+Each training batch contains both real-world samples for unsupervised depth estimation and synthetic color samples.
+Optimizer Step I: The Depth Network is trained in an unsupervised manner on real-world samples using our proposed Auto-UTSF mechanism, while the Color Network is trained in a supervised manner on real-world data to predict color from depth.
+We set $\gamma=10^{-3}$, and $\mathcal{L}^{\text{sm}}$ denotes the smoothness loss.
+The overall loss for this step is defined as
+$\mathcal{L}_{\mathcal{O}_{1}}=\mathcal{L}_{\text{d2c}}+\mathcal{L}_{\text{ud}}+\gamma\mathcal{L}^{\text{sm}}.$ (12)
+Optimizer Step II: The Depth Network is further trained in a supervised manner using the proposed Syn-Real CutMix samples. The corresponding optimization objective is given by
+$\mathcal{L}_{\mathcal{O}_{2}}=\mathcal{L}_{\text{sd}}+\gamma\mathcal{L}^{\text{sm}}.$ (13)
 Figure 11: Visual of Depth Estimation on KITTI: Our method yields clearer depth estimates with higher correlation to corresponding colors, sharper boundaries, and fewer areas with chaotic depth clouds.
@@
 In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.I, Tab.II, and Tab.IV).
-Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI).
-Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.V), perform comparisons at higher resolutions (Tab.III), and benchmark against the lightest and most effective methods (Tab.VIII).
+Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VIII, Tab.VII and Tab.VI).
+Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.V), perform comparisons at higher resolutions (Tab.III), and benchmark against the lightest and most effective methods (Tab.IX).
 At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig.
-13, Fig.
-15, and Fig.
-16).
+11, Fig.
+12, and Fig.
+13).
 IV-A Datasets
-We utilized two unsupervised monocular depth estimation datasets: KITTI[25] and Cityscapes[11], along with synthetic datasets vKitti2[12], SHIFT[23], and UrbanSyn[24].
+We utilized two unsupervised monocular depth estimation datasets: KITTI[10] and Cityscapes[7], along with synthetic datasets vKITTI2[4], SHIFT[51], and UrbanSyn[gómez2023one].
 KITTI contains 39,810 training triplets, 4,424 for validation, and 697 for evaluation, featuring diverse scenes with rich colors and occasional non-rigid objects.
@@
 We demonstrate the advantages of using MixSyn in Tab.
-VII.
+VIII.
 TABLE I: Experiments on the KITTI Eigen Split using only the Monocular Sequence dataset for training.
-Method Sup. Size Error Metric Accuracy Metric
+Method Sup. [^1] Sup.(Supervision) in the tables means what trainning set is used, Video(m),Stereo(s). Size Error Metric Accuracy Metric
 abs_rel sq_rel rmse rmse_log a1 a2 a3
-Monodepth2[9] m 192x640 0.115 0.903 4.863 0.193 0.877 0.959 0.981
-CADepth[29] m 192x640 0.105 0.769 4.535 0.181 0.892 0.964 0.983
-DIFFNet[31] m 192x640 0.102 0.749 4.445 0.179 0.897 0.965 0.983
-MonoFormer[60] m 192x640 0.106 0.839 4.627 0.183 0.889 0.962 0.983
-Lite-Mono[1] m 192x640 0.101 0.729 4.454 0.178 0.897 0.965 0.983
-MonoViT[2] m 192x640 0.099 0.708 4.372 0.175 0.900 0.967 0.984
-Back2Color m 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
-DIFFNet[31] m 320x1024 0.097 0.722 4.345 0.174 0.907 0.967 0.984
-Lite-[1] m 320x1024 0.097 0.710 4.309 0.174 0.905 0.967 0.984
-MonoViT[2] m 320x1024 0.096 0.714 4.292 0.172 0.908 0.968 0.984
-Back2Color m 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 IV-B Implementation Details
-Our training is implemented by Pytorch[61] with NVIDIA RTX 3090Ti(24GB).
-We have used ADAM[62] as the optimizer.
-We have set the learning rate to $1\times 10^{-4}$, the same as in Monodepth2[9], and the learning rate will decay to $1\times 10^{-5}$ after 15 epochs.
-The other parameters are the same as $(\beta_{1}$, $\beta_{2})$ = $(0.9,0.999)$.
-We set the SSIM weight to $0.85$ and the smoothness weight to $\alpha=1e-3$.
-Furthermore, we utilized a pretrained segmentation network[52] trained on the Cityscapes semantic segmentation dataset.
-We evaluate the performance using depth estimation metrics[25, 63]: rmse(log), rmse, Sq.relative, Abs relative, and Accuracy.
+All experiments are implemented using PyTorch [45] and conducted on a single NVIDIA RTX 3090Ti GPU with 24 GB memory.
+We adopt the Adam optimizer [31] with an initial learning rate of $1\times 10^{-4}$, following the setting of Monodepth2 [14].
+The learning rate is decayed to $1\times 10^{-5}$ after 15 epochs, while the remaining hyperparameters are set to $(\beta_{1},\beta_{2})=(0.9,0.999)$.
+The SSIM loss weight is set to $0.85$, and the smoothness regularization weight is set to $\alpha=1\times 10^{-3}$.
+Depth estimation performance is evaluated using standard metrics [10], including $\mathrm{AbsRel}$, $\mathrm{SqRel}$, $\mathrm{RMSE}$, $\mathrm{RMSE}_{\log}$, and accuracy thresholds.
+To ensure computational efficiency and real-time applicability in urban driving scenarios, we adopt a lightweight CNN-based backbone for feature extraction, specifically the Visual Attention Network (VAN) [19], which captures long-range dependencies with low computational cost.
+A minor architectural adaptation is introduced to preserve higher-resolution features for dense depth prediction, while keeping the original backbone structure unchanged.
+The extracted multi-level features are subsequently processed by a high-resolution depth decoder [22] to recover fine-grained spatial details and generate depth maps at the input resolution.
+Experimental results demonstrate that the proposed framework consistently outperforms commonly used CNN-based backbones, such as ResNet-34, HRDepth, and DIFFNet, and achieves competitive performance compared to Transformer-based methods [63, 62] with significantly lower computational complexity.
+VAN0-VAN3 denote four variants of the VAN backbone with different parameter scales, as summarized in TableIX.
+Although VAN is adopted as the default backbone for efficiency, Experimental results demonstrate that the proposed framework consistently improves performance when integrated with commonly used CNN-based backbones, such as ResNet-34, HRDepth, and DIFFNet, and achieves competitive performance compared to Transformer-based methods [63, 62] with significantly lower computational complexity, as shown in Fig.IX.
 TABLE II: Experiments on the KITTI Eigen Split using only the Monocular Sequence and Stereo datasets for training.
-Method Sup. Size Error Metric Accuracy Metric
+Method Sup.11footnotemark: 1 Size Error Metric Accuracy Metric
 abs_rel sq_rel rmse rmse_log a1 a2 a3
-Monodepth2[9] ms 192x640 0.106 0.818 4.750 0.196 0.874 0.957 0.979
-HR-Depth[30] ms 192x640 0.107 0.785 4.612 0.185 0.887 0.962 0.982
-CADepth[29] ms 192x640 0.102 0.752 4.504 0.181 0.894 0.964 0.983
-DIFFNet[31] ms 192x640 0.101 0.749 4.445 0.179 0.898 0.965 0.983
-MonoViT[2] ms 192x640 0.098 0.683 4.333 0.174 0.904 0.967 0.984
-Back2Color ms 192x640 0.092 0.685 4.229 0.170 0.911 0.968 0.984
-Monodepth2[9] ms 320x1024 0.106 0.806 4.630 0.193 0.876 0.958 0.980
-HR-Depth[30] ms 320x1024 0.101 0.716 4.395 0.179 0.894 0.966 0.983
-CADepth[29] ms 320x1024 0.096 0.694 4.264 0.173 0.908 0.968 0.984
-DIFFNet[31] ms 320x1024 0.094 0.678 4.250 0.172 0.911 0.968 0.984
-MonoViT[2] ms 320x1024 0.093 0.671 4.202 0.169 0.912 0.969 0.984
-Back2Color ms 320x1024 0.087 0.644 4.079 0.166 0.920 0.970 0.985
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 TABLE III: Experiments on the KITTI Eigen Split at 384 $\times$ 1280 resolution.
-Method Sup. Error Metric Accuracy Metric
+Method Sup.11footnotemark: 1 Error Metric Accuracy Metric
 abs_rel sq_rel rmse rmse_log a1 a2 a3
@@
 The evaluation results are summarized in Tab.I and Tab.II with KITTI Eigen Split[10] and Tab.IV with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.11 and Fig.12.
-Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[1, 2] with larger params and more complex structures(Tab.VIII).
-Figure 14: Temporal-Spatial Uncertainty Fusion: White represents higher reliability in spatial reprojection, while black represents higher reliability in temporal reprojection, as shown in the legend in the bottom right corner.
-We can observe that non-rigid moving objects in the image tend to reproject more in space (white).
-Meanwhile, occluded areas and distant regions tend to reproject more in the temporal sequence (black), and the remaining areas show no significant tendency (gray).
-TABLE IV: Experiments on the Cityscapes benchmark with 192$\times$640. Methods* requiring semantic information and Back2Color requiring stereo set.
+Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[62, 63] with larger params and more complex structures(Tab.IX).
+[…]
 Method Error Metric Accuracy Metric
@@
 RM-Depth[27] 0.090 0.825 5.503 0.143 0.913 0.980 0.993
-Back2color* 0.076 0.784 5.159 0.127 0.937 0.984 0.994
+Back2color 0.076 0.784 5.159 0.127 0.937 0.984 0.994
 TABLE V: Experiments with methods using the Synthetic dataset (vKITTI2) at 192×640 resolution and training with monocular sequential only.
@@
 abs_rel sq_rel rmse rmse_log a1 a2 a3
-MonoDEVSNet[18] 192x640 0.102 0.685 4.303 0.178 0.894 0.966 0.984
-Back2Color 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
-Swami et al.[43] 320x1024 0.103 0.654 4.300 0.178 0.891 0.966 0.984
-Back2Color 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
+[…]
+[…]
+[…]
+[…]
 IV-D Ablation Study
+TABLE VI: Ablation experiments to test different modules. The experiments use a stereo setup(ms) with a resolution of 320×1024. Row 1 shows the results of Monodepth2[14] as a reference.
+Modules Error Metric Accuracy Metric
+abs_rel sq_rel rmse rmse_log a1 a2 a3
+Backbone 0.094 0.703 4.435 0.173 0.909 0.969 0.984
+Backbone+MixSyn 0.095 0.724 4.672 0.172 0.910 0.969 0.984
+Backbone+MixSyn+B2C 0.090 0.688 4.320 0.169 0.915 0.970 0.985
+Backbone+MixSyn+B2C+UTSF 0.087 0.644 4.079 0.166 0.920 0.970 0.985
 As illustrated in Tab.
-VI, we conducted experiments on our Back2Color framework by replacing different model structures to demonstrate the effectiveness of our method.
-While the Depth Net learned to predict dense depth from synthetic data, it also disrupted the model’s latent memory in unsupervised learning and learned to generate sharp boundaries of obstacles.
-Based on the metrics in Tab.
-VI, there is an improvement across all metrics, particularly the $a1$ and $rmse\_log$ metrics representing detail accuracy.
-At the same time, we also observed that there was no significant difference between only using SHIFT and vKitti2.
-In Tab.
-VI, we also tested training VADepth from scratch without ImageNet pre-training.
-The results demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation.
+VII, we conducted experiments on our Back2Color framework by replacing different network structures to demonstrate the effectiveness of our new framework.
+The results also demonstrated that using a pre-trained checkpoint enhances the network’s ability to process natural images, significantly benefiting depth estimation.
+Next, we conducted ablation experiments to separately evaluate each of the proposed modules, ad shown in Tab.VI including the MixSyn dataset, the Back2Color training framework, and the auto-learning uncertainty temporal-spatial fusion method (UTSF).
+The results demonstrated that simply combining real and synthetic data for training did not improve performance.
+However, applying the Back2Color framework for color domain adaptation effectively addressed this issue.
+Additionally, Auto-UTSF further improved performance by leveraging advantageous information from both the temporal and spatial domains.
 Then we compared our method with other noval approaches also joint training with synthetic datasets. Our improvement remains significant in comparison, as shown in Tab.V.
 Figure 12: The visual results on Cityscapes: Our method yields clearer depth estimates with higher correlation to corresponding colors, sharper boundaries, and fewer areas with chaotic depth clouds.
-TABLE VI: Experiments on the Kitti Eigen Split at 192×640 resolution with different models for Back2Color.VAN* means not apply ImageNet pretrained checkpoints.
+TABLE VII: Experiments on the KITTI Eigen split at 192×640 resolution.
+Results show that Back2Color consistently improves performance across different backbones, indicating that the gain is not solely attributed to backbone selection.
+VAN* denotes training without ImageNet pretraining.
 Method Model Error Metric Accuracy Metric
@@
 Back2Color VAN2 0.092 0.685 4.229 0.170 0.911 0.968 0.984
-Additionally, we used various synthetic datasets such as vKitti2, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Tab.VII.
+Additionally, we used various synthetic datasets such as vKITTI2, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Tab.VIII.
 The results consistently achieved excellent improvement, demonstrating the strong adaptability of our method.
@@
 Meanwhile, our VAN2 model, which has a parameter count similar to MonoViT, shows a significant performance improvement, making it the current state-of-the-art (SOTA) model.
-TABLE VII: Ablation Study on the impact of Various Synthetic Datasets.
+TABLE VIII: Ablation Study on the impact of Various Synthetic Datasets.
 Size Syn Number of Samples Error Metric Accuracy Metric
 KITTI vKITTI2 SHIFT UrbanSyn abs_rel sq_rel rmse rmse_log a1 a2 a3
-192x640 vKitti2 39810 21260 0 0 0.094 0.694 4.296 0.173 0.907 0.968 0.984
-192x640 SHIFT 39810 0 153000 0 0.092 0.676 4.238 0.171 0.909 0.969 0.984
-192x640 Urban 39810 0 0 7539 0.093 0.644 4.198 0.172 0.907 0.968 0.984
-192x640 MixSyn 39810 21260 153000 7539 0.092 0.685 4.229 0.170 0.911 0.968 0.984
-320x1024 vKitti2 39810 21260 0 0 0.090 0.677 4.173 0.169 0.914 0.969 0.984
-320x1024 vKitti2 31848 7962 0 0 0.090 0.675 4.143 0.168 0.917 0.970 0.985
-320x1024 vKitti2 19905 19905 0 0 0.091 0.684 4.192 0.169 0.913 0.969 0.985
-320x1024 SHIFT 39810 0 153000 0 0.089 0.651 4.158 0.167 0.916 0.970 0.985
-320x1024 SHIFT 31848 0 7962 0 0.090 0.684 4.158 0.168 0.916 0.970 0.984
-320x1024 SHIFT 19905 0 19905 0 0.090 0.680 4.144 0.167 0.915 0.970 0.985
-320x1024 Urban 39810 0 0 7539 0.090 0.654 4.084 0.167 0.915 0.970 0.985
-320x1024 MixSyn 39810 21260 153000 7539 0.087 0.644 4.079 0.166 0.920 0.970 0.985
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Figure 13: The visual results of joint training with different synthetic datasets.
 The model trained with SHIFT jointly learns more features of pedestrians while treating glass as transparent entities.
-TABLE VIII: Ablation study on models with params and computational complexities(GFLOPs).
+TABLE IX: Ablation study on models with params and computational complexities(GFLOPs).
 Method Model Size Params(M) Flops(G) Error Metric Accuracy Metric
 abs_rel sq_rel rmse rmse_log a1 a2 a3
-MonoViT[2] small 320x1024 33.63 159.14 0.093 0.671 4.202 0.169 0.912 0.969 0.985
-Lite-mono[1] 8M 320x1024 8.76 29.88 0.097 0.710 4.309 0.174 0.905 0.967 0.984
-Back2Color VAN0 320x1024 5.17 17.64 0.094 0.706 4.257 0.171 0.910 0.969 0.984
-Back2Color VAN1 320x1024 18.61 51.48 0.090 0.664 4.139 0.168 0.915 0.970 0.984
-Back2Color VAN2 320x1024 31.32 76.01 0.087 0.644 4.079 0.166 0.920 0.970 0.985
-Back2Color VAN3 320x1024 49.5 114.63 0.087 0.653 4.068 0.165 0.919 0.971 0.985
-Back2Color VAN0 384x1280 5.17 17.64 0.093 0.676 4.169 0.171 0.911 0.969 0.984
-Back2Color VAN1 384x1280 18.61 51.48 0.090 0.664 4.102 0.168 0.917 0.970 0.984
-Back2Color VAN2 384x1280 31.32 76.01 0.086 0.629 4.054 0.165 0.920 0.970 0.985
-Back2Color VAN3 384x1280 49.5 114.63 0.087 0.652 4.062 0.165 0.920 0.970 0.984
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 IV-E Visualization Analysis
@@
 Firstly, it’s undeniable that our method, through joint training with synthetic datasets, achieves clearer boundaries and more detailed depth estimations compared to training solely with unsupervised depth estimation, as depicted in Fig.
-13.
+11.
 The model successfully captures features of obstacles that unsupervised learning alone struggles to discern.
@@
 IV-F Discussion
-IV-F1 Why can the network transform information beyond depth
+Why can the network transform information beyond depth?
 In Fig.
@@
 To validate, we directly trained the network to predict color images from synthesized depth in the synthetic dataset (Fig.
-17).
+14).
 The absence of lane markings and shadows in the restored color images confirms our hypothesis.
-IV-F2 Why learn Depth and Color Net simultaneously online
-the primary objective of our method is depth prediction.
-As the Depth Net undergoes continuous updates to enhance its accuracy and performance, it is imperative to ensure that the Color Net also adapts accordingly.
-This synchronization is crucial for achieving superior color transformation effects.
-By reprojecting the Color Net’s adjustments with the updates in the Depth Net, we aim to optimize the overall performance of the model.
-This dynamic adjustment process ensures that the color transformation remains effective and reproject with the evolving depth predictions.
-IV-F3 Can the VAN-based encoder replace the ResNet structure in the Pose encoder
-so far, we have not observed any new research that can replace the ResNet-based PoseEncoder[9], whether using CNN-based or Transformer-based methods.
-We have attempted some alternatives, but none have improved depth estimation performance; in fact, many have resulted in performance degradation.
-In this study, we also tried using a VAN-based model to replace the ResNet-based PoseEncoder, but still did not observe any significant improvement.
-We believe that the pose estimation task is a small parameter regression task that is relatively simple compared to depth prediction and does not require a powerful network.
+In practice, the ambiguity is significantly reduced by the strong appearance consistency within a target domain (e.g., KITTI or Cityscapes), enabling the depth-to-color network to learn a stable, domain-specific mapping.
 V Conclusion
-In this paper, we presented Back to the Color(Back2Color), a novel framework designed to enhance unsupervised depth estimation by leveraging synthetic data.
-Our method addresses the challenge of aligning synthetic world with real world colors, enabling the transformation of predicted synthetic depth into realistic colors with high fidelity.
-By joint training on synthetic and real-world datasets using Back2Color, we observed substantial improvements in depth estimation performance.
-Additionally, we designed a depth net based on the Visual Attention Network (VAN), which allowed our CNN-based network to surpass several Transformer-based depth models.
-Furthermore, we improved the uncertainty-based loss fusion method by combining the advantages of temporal and spatial reprojection, thoroughly addressing issues related to non-rigid motion.
-Experimental results demonstrated the effectiveness of our method on both the KITTI and Cityscapes datasets.
-Overall, Back2Color offers a promising solution for advancing unsupervised depth estimation by effectively integrating synthetic data into the training process.
+This paper presented Back2Color, an unsupervised monocular depth estimation framework for robust depth perception in urban driving scenarios.
+By learning a depth-to-color mapping from real-world data and transferring it to synthetic scenes, the proposed method effectively reduces the synthetic-to-real appearance gap and improves depth estimation under real traffic conditions.
+In addition, an uncertainty-aware temporal-spatial fusion strategy is introduced to handle non-rigid motion and occlusions, leading to more reliable depth predictions in dynamic environments.
+Experimental results on standard driving benchmarks demonstrate consistent performance improvements with favorable computational efficiency, highlighting the practicality of the proposed approach for vision-based perception in intelligent transportation systems.